### v3

# Differentially Private Bilevel Optimization

Abstract
We present differentially private (DP) algorithms for bilevel optimization, a problem class that received significant attention lately in various machine learning applications. These are the first algorithms for such problems under standard DP constraints, and are also the first to avoid Hessian computations which are prohibitive in large-scale settings. Under the well-studied setting in which the upper-level is not necessarily convex and the lower-level problem is strongly-convex, our proposed gradient-based $(\epsilon,\delta)$-DP algorithm returns a point with hypergradient norm at most $\widetilde{\mathcal{O}}\left((\sqrt{d_{\mathrm{up}}}/\epsilon n)^{1/2}+(\sqrt{d_{\mathrm{low}}}/\epsilon n)^{1/3}\right)$ where $n$ is the dataset size, and $d_{\mathrm{up}}/d_{\mathrm{low}}$ are the upper/lower level dimensions. Our analysis covers constrained and unconstrained problems alike, accounts for mini-batch gradients, and applies to both empirical and population losses. As an application, we specialize our analysis to derive a simple private rule for tuning a regularization hyperparameter.
1 Introduction
Bilevel optimization (BO) is a fundamental framework for solving optimization objectives of hierarchical structure, in which constraints are defined themselves by an auxiliary optimization problem. Formally, it is defined as
$\mathrm{minimize}_{\mathbf{x}\in\mathcal{X}}$~~ $F(\mathbf{x}):=f(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))$~~~~~ (BO)
$\mathrm{subject~to}$~~~ $\mathbf{y}^{*}(\mathbf{x})\in\arg\min_{\mathbf{y}}g(\mathbf{x},\mathbf{y})~,$
where $F:\mathbb{R}^{d_{x}}\to\mathbb{R}$ is referred to as the hyperobjective, $f:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$ as the upper-level (or outer) objective, and $g:\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}\to\mathbb{R}$ as the lower-level (or inner) objective. While BO is well studied for over half a century (Bracken and McGill, 1973), it has recently received significant attention due to its diverse applications in machine learning (ML). These include hyperparameter tuning (Bengio, 2000; Maclaurin et al., 2015; Franceschi et al., 2017, 2018; Lorraine et al., 2020; Engstrom et al., 2025), meta-learning (Andrychowicz et al., 2016; Bertinetto et al., 2018; Rajeswaran et al., 2019; Ji et al., 2020), neural architecture search (Liu et al., 2018), invariant learning (Arjovsky et al., 2019; Jiang and Veitch, 2022), and data reweighting (Grangier et al., 2023; Fan et al., 2024; Pan et al., 2025). In these applications, both the upper and lower level objectives typically represent some loss over data, and are given by empirical risk minimization (ERM) problems over a dataset $S=\{\xi_{1},\dots,\xi_{n}\}\in\Xi^{n}:$ [^1] It is possible for the datasets with respect to $f$ and $g$ to be distinct (e.g., validation/training), perhaps of different sizes. In this case, $S$ is the entire dataset, and $n$ is the total number of samples. Letting $f(\cdot;\xi_{i})=0$ or $g(\cdot;\xi_{i})=0$ for certain indices in order to exclude corresponding data points from either objective, will not affect our results.
$f(\mathbf{x},\mathbf{y}):=f_{S}(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}f(\mathbf{x},\mathbf{y},\xi_{i})~,\qquad g(\mathbf{x},\mathbf{y}):=g_{S}(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}g(\mathbf{x},\mathbf{y},\xi_{i})~,$ (ERM)
where both objectives are often proxies of stochastic (i.e., population) losses with respect to a distribution $\mathcal{P}$ over $\Xi:$
$f(\mathbf{x},\mathbf{y}):=f_{\mathcal{P}}(\mathbf{x},\mathbf{y})=\mathbb{E}_{\xi\sim\mathcal{P}}[f(\mathbf{x},\mathbf{y};\xi)]~,\qquad g(\mathbf{x},\mathbf{y}):=g_{\mathcal{P}}(\mathbf{x},\mathbf{y})=\mathbb{E}_{\xi\sim\mathcal{P}}[g(\mathbf{x},\mathbf{y};\xi)]~.$ (Pop)
In this work, we study bilvel optimization under differential privacy (DP) (Dwork et al., 2006). As ML models are deployed in an ever-growing number of applications, protecting the privacy of the data on which they are trained is a major concern, and DP has become the gold-standard for privacy preserving ML (Abadi et al., 2016). Accordingly, DP optimization is extensively studied, with a vast literature focusing both on empirical and stochastic objectives under various settings (Chaudhuri et al., 2011; Kifer et al., 2012; Bassily et al., 2014; Wang et al., 2017; Bassily et al., 2019; Wang et al., 2019; Feldman et al., 2020; Tran and Cutkosky, 2022; Gopi et al., 2022; Arora et al., 2023; Carmon et al., 2023; Ganesh et al., 2024; Lowy et al., 2024).
Nonetheless, private algorithms for BO remain largely underexplored, and in particular, no gradient-based algorithm (aka first-order, which uses only gradient queries) that solves BO problems under DP, is known to date. This is no coincidence: until recently, no first-order methods with finite time guarantees were known even for non-private bilevel problems. This follows the fact (Ghadimi and Wang, 2018, Lemma 2.1) that under mild regularity assumptions, the so-called hypergradient takes the form:
$\nabla F(\mathbf{x})=\nabla_{x}f(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))-\nabla^{2}_{xy}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))[\nabla^{2}_{yy}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))]^{-1}\nabla_{y}f(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))~.$ (1)
Consequently, directly applying a “gradient” method to $F$ requires inverting Hessians of the lower-level problem at each time step, thus limiting applicability in contemporary high-dimensional applications. Following various approaches to tackle this challenge (see Section 1.2), recent breakthroughs were finally able to provide fully first-order methods for BO with non-asymptotic guarantees (Liu et al., 2022; Kwon et al., 2023; Yang et al., 2023; Chen et al., 2024). These recent algorithmic advancements show promising empirical results in large scale applications, even up to the LLM scale of ${\sim}10^{9}$ parameters (Pan et al., 2025).
On the downside, as we will show, first-order bilevel techniques can lead to major privacy violations. Intuitively, this is due to possibly private information leaking between the inner and outer objectives, since $\nabla F(\mathbf{x})$ itself depends on $\mathbf{y}^{*}(\mathbf{x})$ (as seen in Eq. (1)). In fact, we show in Section 3 that even a single hypergradient computation can completely break privacy, even when the upper-level objective does not depend on the data at all.
The only prior method we are aware of for DP BO was proposed by Chen and Wang (2024), which falls short in two main aspects. First, their algorithm only provides some privacy guarantee which cannot be controlled by the user. Moreover, it requires evaluating and inverting Hessians at each step, namely it is not first-order, which limits scalability; see Section 1.2 for further discussion.
1.1 Our contributions
We present DP algorithms that solve BO problems whenever the the upper-level is smooth but not necessarily convex, and the lower-level problem is smooth and strongly-convex. These are the first bilevel algorithms of any sort under standard central DP constraints, and in particular, are the first private algorithms to do so using only first-order (i.e., gradient) queries of the upper- and lower-level objectives. Our contributions can be summarized as follows:
  • DP Bilevel ERM Algorithm (Theorem 4.1): We present a $(\epsilon,\delta)$-DP first-order algorithm for the bilevel ERM problem (BO/ERM) that outputs with high probability a point with hypergradient norm bounded by
    $\left\|\nabla F_{S}(\mathbf{x})\right\|=\widetilde{\mathcal{O}}\bigg(\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}\bigg)~.$
    Moreover, our algorithm is also applicable to the case where $\mathcal{X}\subsetneq\mathbb{R}^{d_{x}}$ is a non-trivial constraint set, which is common in applications of BO. [^2] For instance, in data reweighting $\mathcal{X}$ is the probability simplex, and in hyperparameter tuning it is the hyperparameter space, which is typically constrained (e.g., non-negative reals). In the constrained setting, we obtain the same guarantee as above in terms of the projected hypergradient (see Section 2 for details).
  • Mini-batch DP Bilevel ERM Algorithm (Theorem 5.1): Aiming for a more practical algorithm, we design a variant of our previous algorithm that relies on mini-batch gradients. For the bilevel ERM problem (BO/ERM), given any batch sizes $b_{\mathrm{in}},b_{\mathrm{out}}\in\{1,\dots,n\}$ for sampling gradients of the inner/outer problems respectively, our algorithm ensures $(\epsilon,\delta)$-DP and outputs with high probability a point with hypergradient norm bounded by
    $\left\|\nabla F_{S}(\mathbf{x})\right\|=\widetilde{\mathcal{O}}\bigg(\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}+\frac{1}{b_{\mathrm{out}}}\bigg)~.$ (2)
    Notably, Eq. (2) is independent of the inner-batch size, yet depends on the outer-batch size, which coincides with known results for “single”-level constrained nonconvex optimization (Ghadimi et al., 2016) (see Remark 5.2 for further discussion). Our mini-batch algorithm is also applicable in the constrained setting $\mathcal{X}\subsetneq\mathbb{R}^{d_{x}}$ with the same guarantee in terms of projected hypergradient.
  • Population loss guarantees (Theorem 6.1): We further provide guarantees for stochastic objectives. In particular, we show that for the population bilevel problem (BO/Pop), our $(\epsilon,\delta)$-DP algorithm outputs with high probability a point with hypergradient norm bounded by
    $\left\|\nabla F_{\mathcal{P}}(\mathbf{x})\right\|=\widetilde{\mathcal{O}}\bigg(\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}+\Big(\frac{d_{x}}{n}\Big)^{1/2}+\frac{1}{n^{1/4}}\bigg)~,$
    with an additional additive $1/b_{\mathrm{out}}$ factor in the mini-batch setting.
  • Application to private hyperparameter tuning (Section 7): We specialize our algorithmic framework to the classic problem of tuning a regularization hyperparameter, when fitting a regularized statistical model. Our analysis of this problem leads to a simple, privacy preserving update rule for tuning the amount of regularization. As far as we know, this is the first such method which avoids selecting a hyperparameter from a predetermined set of candidates (as studied by previous works), but rather updates the hyperparameter privately on the fly.
1.2 Related work
BO was introduced by Bracken and McGill (1973), and grew into a vast body of work, with classical results focusing on asymptotic guarantees for certain specific problem structures (Anandalingam and White, 1990; Ishizuka and Aiyoshi, 1992; White and Anandalingam, 1993; Vicente et al., 1994; Zhu, 1995; Ye and Zhu, 1997). There exist multiple surveys and books covering various approaches for these problems (Vicente and Calamai, 1994; Dempe, 2002; Colson et al., 2007; Bard, 2013; Sinha et al., 2017).
Ghadimi and Wang (2018) observed Eq. (1) under strong-convexity of the inner problem using the implicit function theorem, asserting that the hypergradient can be computed via inverse Hessians, which requires solving a linear system at each point. Many follow up works built upon this second-order approach with additional techniques such as variance reduction, momentum, Hessian sketches, projection-free updates, or incorporating external constraints (Amini and Yousefian, 2019; Yang et al., 2021; Khanduri et al., 2021; Guo et al., 2021; Ji et al., 2021; Chen et al., 2021; Akhtar et al., 2022; Chen et al., 2022; Tsaknakis et al., 2022; Hong et al., 2023; Jiang et al., 2023; Abolfazli et al., 2023; Merchav and Sabach, 2023; Xu and Zhu, 2023; Cao et al., 2024; Dagréou et al., 2024). Only recently, the groundbreaking result of Liu et al. (2022) proved finite-time convergence guarantees for a fully first-order method which is based on a penalty approach. This result was soon extended to stochastic objectives (Kwon et al., 2023), with the convergence rate later improved by (Yang et al., 2023; Chen et al., 2024, 2025), and also extended to constrained bilevel problems (Yao et al., 2024; Kornowski et al., 2024). The first-order penalty paradigm also shows promise for some bilevel problems in which the inner problem is not strongly-convex (Shen and Chen, 2023; Kwon et al., 2024; Lu and Mei, 2024), which is generally a highly challenging setting (Chen et al., 2024; Bolte et al., 2025). Moreover, Pan et al. (2025) provided an efficient implementation of this paradigm, showing its effectiveness for large scale applications.
As to DP optimization, there is an extensive literature both for ERM and stochastic losses, for either convex objectives (Chaudhuri et al., 2011; Kifer et al., 2012; Bassily et al., 2014; Wang et al., 2017; Bassily et al., 2019; Feldman et al., 2020; Gopi et al., 2022; Carmon et al., 2023) as well as for nonconvex objectives (Wang et al., 2019; Tran and Cutkosky, 2022; Arora et al., 2023; Ganesh et al., 2024; Lowy et al., 2024; Zhang et al., 2024; Kornowski et al., 2025).
To the best of our knowledge, the only existing result for DP BO prior to our work is the result of Chen and Wang (2024), which differs than ours in several aspect. Their proposed algorithm is second-order, requiring evaluating Hessians and inverting them at each time step, which we avoid altogether. Moreover, Chen and Wang (2024) study the local DP model (Kasiviswanathan et al., 2011), in which each user (i.e., $\xi_{i}$) does not reveal its individual information. Due to this more challenging setting, they only derive guarantees for some finite privacy budget $\epsilon<\infty$, even as the dataset size grows. We study the common central DP model, in which a trusted curator acts on the collected data and releases a private solution, and we are able to provide any desired privacy and accuracy guarantees with sufficiently many samples. Our work is the first to study BO in this common DP setting. We also note that Fioretto et al. (2021) studied the related problem of DP in Stackelberg games, which are certain bilevel programs which arise in game theory, aiming at designing coordination mechanisms that maintain each agent’s privacy.
After the initial version of this work became available on arXiv (Kornowski, 2024), Lowy and Liu (2025) further studied DP BO, proposing second-order algorithms which at the computational cost of inverting Hessians, are able to achieve gradient bounds which, interestingly, remove the dependence on the lower-level dimension $d_{y}$ presented in this work.
2 Preliminaries
Notation and terminology.
[…]
Differential privacy.
Two datasets $S,S^{\prime}\in\Xi^{n}$ are said to be neighboring if they differ by only one data point. A randomized algorithm $\mathcal{A}:\Xi^{n}\to\mathcal{R}$ is called $(\epsilon,\delta)$ differentially private (or $(\epsilon,\delta)$-DP) for $\epsilon,\delta>0$ if for any two neighboring datasets $S\sim S^{\prime}$ and measurable $E\subseteq\mathcal{R}$ in the algorithm’s range: $\Pr[\mathcal{A}(S)\in E]\leq e^{\epsilon}\Pr[\mathcal{A}(S^{\prime})\in E]+\delta$ (Dwork et al., 2006). In Appendix A we recall some well known DP basics used in our analyses: composition and advanced composition, the Gaussian mechanism, and privacy amplification by subsampling.
Gradient mapping.
Given a point $\mathbf{x}\in\mathbb{R}^{d}$, and some $\mathbf{v}\in\mathbb{R}^{d},~\eta>0$, we denote $\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x}):=\frac{1}{\eta}\left(\mathbf{x}-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right),~\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x}):=\arg\min_{\mathbf{u}\in\mathcal{X}}\big[\left\langle\mathbf{v},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}\big]$. In particular, given an $L$-smooth function $F:\mathbb{R}^{d}\to\mathbb{R}$ and $\eta\leq\frac{1}{2L}$, we denote the projected gradient (also known as reduced gradient) and the gradient (or proximal) mapping, respectively, as
$\mathcal{G}_{F,\eta}(\mathbf{x}):=\frac{1}{\eta}\left(\mathbf{x}-\mathcal{P}_{\nabla F,\eta}(\mathbf{x})\right)~,~~~~~\mathcal{P}_{\nabla F,\eta}(\mathbf{x}):=\arg\min_{\mathbf{u}\in\mathcal{X}}\Big[\left\langle\nabla F(\mathbf{x}),\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}\Big]~.$
The projected gradient $\mathcal{G}_{F,\eta}(\mathbf{x})$ generalizes the gradient to the possibly constrained setting: for points $\mathbf{x}\in\mathcal{X}$ sufficiently far from the boundary of $\mathcal{X},~\mathcal{G}_{F,\eta}(\mathbf{x})=\nabla F(\mathbf{x})$, namely it simply reduces to the gradient; see the textbooks (Nesterov, 2013; Lan, 2020) for additional details.
2.1 Setting
We impose the following assumptions which are standard in the BO literature.
Assumption 2.1.
For (BO) with either (ERM) or (Pop), we assume the following hold:
  1. i.
    $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ is a closed convex set;
  2. ii.
    $F(\mathbf{x}_{0})-\inf_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\Delta_{F}$ for some initial point $\mathbf{x}_{0}\in\mathcal{X}$;
  3. iii.
    $f$ is twice differentiable, and $L_{1}^{f}$-smooth;
  4. iv.
    For all $\xi\in\Xi:~f(\cdot,\cdot\,;\xi)$ is $L_{0}^{f}$-Lipschitz (hence, so is $f$);
  5. v.
    $g$ is $L_{2}^{g}$-Hessian-smooth, and for all $\mathbf{x}\in\mathcal{X}:~g(\mathbf{x},\cdot)$ is $\mu_{g}$-strongly-convex;
  6. vi.
    For all $\xi\in\Xi:~g(\cdot,\cdot\,;\xi)$ is $L_{1}^{g}$-smooth (hence, so is $g$).
As mentioned, these assumptions are standard in the study of BO problems and are shared by nearly all of the previously discussed works. In particular, the strong convexity of $g(\mathbf{x},\cdot)$ ensures that $\mathbf{y}^{*}(\mathbf{x})$ is uniquely defined, which is generally required in establishing the regularity of the hyperobjective. Indeed, it is known that dropping this assumption, can, in general, lead to pathological behaviors not amenable for algorithmic guarantees (cf. Chen et al. 2024; Bolte et al. 2025 and discussions therein). For the purpose of differential privacy though, the strong convexity of $g(\mathbf{x},\cdot)$ raises a subtle issue. As the standard assumption in the DP optimization literature is that the component functions are Lipschitz, which allows privatizing gradients using sensitivity arguments, strongly-convex objectives cannot be Lipschitz over the entire Euclidean space. [^3] If $g(\mathbf{x},\cdot\,;\xi)$ were Lipschitz over $\mathbb{R}^{d_{y}}$ for all $\xi\in\Xi$, then so would $g(\mathbf{x},\cdot)$, contradicting strong convexity. Therefore, strongly-convex objectives are regularly analyzed in the DP setting under the additional assumption that the domain of interest is bounded. For bilevel problems, the domain of interest for $\mathbf{y}$ is the lower-level solution set, thus we impose the following assumption.
Assumption 2.2.
There exists a compact set $\mathcal{Y}\subset\mathbb{R}^{d_{y}}$ with $\{\mathbf{y}^{*}(\mathbf{x}):\mathbf{x}\in\mathcal{X}\}\subseteq\mathcal{Y}$, such that for all $\mathbf{x}\in\mathcal{X},\,\xi\in\Xi:~g(\mathbf{x},\cdot\,;\xi)$ is $L_{0}^{g}$-Lipschitz over $\mathcal{Y}$.
Remark 2.3.
Note that $\mathrm{diam}(\mathcal{Y})\leq{2L_{0}^{g}}/{\mu_{g}}$: Indeed, fixing some $\mathbf{x}\in\mathcal{X}$, $g(\mathbf{x},\cdot\,;\xi)$ is $L_{0}^{g}$-Lipschitz over $\mathcal{Y}$ for all $\xi\in\Xi$, thus so is $g(\mathbf{x},\cdot)$. By $\mu_{g}$-strong-convexity, we get for all $\mathbf{y}\in\mathcal{Y}:~\mu_{g}\left\|\mathbf{y}-\mathbf{y}^{*}(\mathbf{x})\right\|$ $\leq\left\|\nabla_{y}g(\mathbf{x},\mathbf{y})\right\|\leq L_{0}^{g}$. Hence $\mathcal{Y}\subseteq\mathbb{B}(\mathbf{y}^{*}(\mathbf{x}),{L_{0}^{g}}/{\mu_{g}})$, which is of diameter ${2L_{0}^{g}}/{\mu_{g}}$.
Following Assumptions 2.1 and 2.2, we denote $\ell:=\max\{L_{0}^{f},L_{1}^{f},L_{0}^{g},L_{1}^{g},L_{2}^{g}\},~\kappa:=\ell/\mu_{g}$.
3 Warm up: Privacy can leak from lower to upper level
In this section, we provide an illuminating example of privacy breaking in first-order BO, which motivates the algorithmic approach presented later in the paper. In particular, we exemplify that even when the upper-level problem supposedly does not depend on the dataset at all, the bilevel structure can give rise to hyper-gradients that are non-private.
Example 3.1.
Given any dataset $S=\{\xi_{1},\dots,\xi_{n}\}\subset\mathbb{R}^{d}$, consider the BO problem corresponding to the objectives over $\mathbb{R}^{2d}:$
$f(\mathbf{x},\mathbf{y})=\frac{1}{2}\left\|\mathbf{x}+\mathbf{y}\right\|^{2}~,\qquad g(\mathbf{x},\mathbf{y})=\frac{1}{2}\sum_{i=1}^{n}\left\|\mathbf{y}-\xi_{i}\right\|^{2}~.$
These are smooth non-negative objectives, and $g$ is strongly-convex with respect to $\mathbf{y}$, and so we see that Assumptions 2.1 and 2.2 hold. Further note that the upper-level objective $f$ does not depend on the data $S$ at all. Nonetheless, we will show that a single hypergradient computation can break privacy.
To see this, note that for any $\mathbf{x}\in\mathbb{R}^{d}:~\mathbf{y}^{*}(\mathbf{x})=\arg\min_{\mathbf{y}}g(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$ since the squared loss is minimized by the average. Furthermore, $\nabla_{x}f(\mathbf{x},\mathbf{y})=\mathbf{x}+\mathbf{y}$ and $\nabla^{2}_{xy}g\equiv\mathbf{0}$, so by Eq. (1) we get that $\nabla F(\mathbf{x})=\mathbf{x}+\mathbf{y}^{*}(\mathbf{x})=\mathbf{x}+\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$. In particular, computing the hypergradient at the origin simply returns the dataset’s mean $\nabla F(\mathbf{0})=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$, which is well-known to be a non-private statistic of the dataset $S$ (cf. Kamath et al., 2019).
The algorithms to follow incorporate a fix to overcome such privacy leaks: instead of ever computing $\mathbf{y}^{*}(\mathbf{x})$ (as non-private algorithms do up to exponentially small precision), we estimate it using an auxiliary private method, and use the resulting private point $\widetilde{\mathbf{y}}^{*}(\mathbf{x})$ to estimate the hypergradient at $\mathbf{x}$. [^4] To be precise, in order to also avoid Hessian computations, $\nabla F(\mathbf{x})$ will be estimated by the difference of gradients at two private points close to $\mathbf{y}^{*}(\mathbf{x})$. This leads to non-negligible bias in the hypergradient estimators compared to non-private first-order BO, that we need to account for in our analysis. We further discuss this in Section 4.1.
4 Algorithm for DP bilevel ERM
In this section, we consider the bilevel ERM problem (BO/ERM), for which we denote the hyperobjective by $F_{S}$. Our algorithm is presented in Algorithm 1. We prove the following result:
Theorem 4.1.
Assume 2.1 and 2.2 hold, and that $\alpha\leq\ell\kappa^{3}\min\{\frac{1}{2\kappa},\frac{L_{0}^{g}}{L_{0}^{f}},\frac{L_{1}^{g}}{L_{1}^{f}},\frac{\Delta_{F}}{\ell\kappa}\}$. Then there is a parameter assignment $\lambda\asymp\ell\kappa^{3}\alpha^{-1},~\sigma^{2}\asymp\ell^{2}\kappa^{2}T\log(T/\delta)\epsilon^{-2}n^{-2},~\eta\asymp\ell^{-1}\kappa^{-3},~T\asymp\Delta_{F}\ell\kappa^{3}\alpha^{-2}$, such that running Algorithm 1 satisfies $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
$\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha=\widetilde{\mathcal{O}}\bigg(K_{1}\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+K_{2}\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}\bigg)~,$
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{3/4}\kappa^{5/4}),~K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{5/6}\kappa^{11/6})$.
Remark 4.2.
Recall that when $\mathcal{X}=\mathbb{R}^{d_{x}}$, then $\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})=\nabla F_{S}(\mathbf{x}_{\mathrm{out}})$.
Remark 4.3.
Algorithm 1 generalizes previously studied algorithms in two special cases. First, in the non-private case (when $\epsilon=\infty$ or $\delta=1$) the algorithm can skip the private inner loop and simply solve the inner problems instead in $\widetilde{\mathcal{O}}(1)$ steps, thus reducing to the first-order non-private bilevel algorithms analyzed by (Kwon et al., 2023; Chen et al., 2024, 2025). Second, if the lower-level objective is independent of $\mathbf{x}$, the partial derivative $\nabla_{x}g\equiv\mathbf{0}$ vanishes and ${\widetilde{\mathbf{y}}}_{t}\approx\arg\min_{\mathbf{y}}g(\mathbf{x}_{t},\mathbf{y})$ remains constant independently of $t$. In this case, Algorithm 1 simply reduces to DP-GD with respect to the first variable of $f$, and indeed, the $\mathcal{O}(({\sqrt{d_{x}}}/{\epsilon n})^{1/2})$ gradient bound we obtain matches the known result of DP-GD for smooth-nonconvex “single-level” optimization (Wang et al., 2017).
1:Input: Initialization $(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$, privacy budget $(\epsilon,\delta)$, penalty $\lambda>0$, noise level $\sigma^{2}>0$, step size $\eta>0$, iteration budget $T\in\mathbb{N}$.
2:for $t=0,\dots,T-1$ do
3:  Apply $(\frac{\epsilon}{\sqrt{18T}},\frac{\delta}{3(T+1)})$-DP-Loc-GD (Algorithm 2) to solve $\triangleright$ Strongly-convex problems
${\widetilde{\mathbf{y}}}_{t}\approx\arg\min_{\mathbf{y}}g(\mathbf{x}_{t},\mathbf{y})~,~~~~~\widetilde{\mathbf{y}}^{\lambda}_{t}\approx\arg\min_{\mathbf{y}}\left[f(\mathbf{x}_{t},\mathbf{y})+\lambda\cdot g(\mathbf{x}_{t},\mathbf{y})\right]$
4:  $\widetilde{\mathbf{g}}_{t}=\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t})\right)+\nu_{t}$, where $\nu_{t}\sim\mathcal{N}(\mathbf{0},\sigma^{2}I_{d_{x}})$
5:  $\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\left\langle\widetilde{\mathbf{g}}_{t},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}_{t}\right\|^{2}\right\}$ $\triangleright$ If $\mathcal{X}=\mathbb{R}^{d_{x}}$, then $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\widetilde{\mathbf{g}}_{t}$
6:end for
[…]
8:Output: $\mathbf{x}_{t_{\mathrm{out}}}$.
Algorithm 1 DP Bilevel
1:Input: Objective $h:\mathbb{R}^{d_{y}}\to\mathbb{R}$, initialization $\mathbf{y}_{0}\in\mathcal{Y}$, privacy budget $(\epsilon^{\prime},\delta^{\prime})$, number of rounds $M\in\mathbb{N}$, noise level $\sigma_{\mathrm{GD}}^{2}>0$, step sizes $(\eta_{t})_{t=0}^{T-1}$, iteration budget $T_{\mathrm{GD}}\in\mathbb{N}$, radii $(R_{m})_{m=0}^{M-1}>0$.
[…]
3:for $m=0,\dots,M-1$ do
4:  for $t=0,\dots,T_{\mathrm{GD}}-1$ do
5:   $\mathbf{y}_{t+1}^{m}=\mathrm{Proj}_{\mathbb{B}(\mathbf{y}_{0}^{m},R_{m})}\left[\mathbf{y}_{t}^{m}-\eta_{t}\left(\nabla h(\mathbf{y}_{t}^{m})+\nu_{t}\right)\right]$, where $\nu_{t}^{m}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathrm{GD}}^{2}I_{d_{y}})$
6:  end for
7:  $\mathbf{y}_{0}^{m+1}=\frac{1}{T}\sum_{t=0}^{T-1}\mathbf{y}_{t}^{m}$
8:end for
9:Output: $\mathbf{y}_{\mathrm{out}}=\mathbf{y}^{M}_{0}$.
Algorithm 2 DP-Loc-GD
4.1 Analysis overview
In this section, we will go over the main ideas that appear in the proof of Theorem 4.1, all of which are provided in full detail in Section 8. We start by introducing some useful notation: Given $\lambda>0$, we denote the penalty function
$\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}):=f(\mathbf{x},\mathbf{y})+\lambda\left[g(\mathbf{x},\mathbf{y})-g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))\right]~,$
$\text{and}~~~\mathcal{L}^{*}_{\lambda}(\mathbf{x}):=\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))~,~~~~\mathbf{y}^{\lambda}(\mathbf{x}):=\arg\min_{\mathbf{y}}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y})~.$
The starting point of our analysis is the following result:
Lemma 4.4 (Kwon et al. 2023; Chen et al. 2024, 2025).
If $\lambda\geq 2L_{1}^{f}/\mu_{g}$, then:  (i) $\left\|\mathcal{L}^{*}_{\lambda}-F\right\|_{\infty}=\mathcal{O}(\ell\kappa/\lambda)$;  (ii) $\left\|\nabla\mathcal{L}^{*}_{\lambda}-\nabla F\right\|_{\infty}=\mathcal{O}(\ell\kappa^{3}/\lambda)$;  (iii) $\mathcal{L}_{\lambda}^{*}$ is $\mathcal{O}(\ell\kappa^{3})$-smooth (independently of $\lambda$).
In other words, the lemma shows that for sufficiently large penalty $\lambda$, $\mathcal{L}_{\lambda}^{*}$ is a smooth approximation of the hyperobjective $F$, and that it suffices to minimize the gradient norm of $\mathcal{L}_{\lambda}^{*}$ in order to get a hypergradient guarantee in terms of $\nabla F$. Moreover, had we computed $\mathbf{y}^{*}(\mathbf{x}),\mathbf{y}^{\lambda}(\mathbf{x})$, we note that $\nabla\mathcal{L}_{\lambda}^{*}$ can be obtained in a first-order fashion, since by construction $\mathcal{L}_{\lambda}^{*}(\mathbf{x})=\arg\min_{\mathbf{y}}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y})$, and therefore by first-order optimality:
$\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x})$ $=\nabla_{x}\mathcal{L}^{*}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))+\nabla_{x}\mathbf{y}^{\lambda}(\mathbf{x})^{\top}\underset{=\mathbf{0}}{\underbrace{\nabla_{y}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))}}$
$=\nabla_{x}f(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))+\lambda\left(\nabla_{x}g(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))-\nabla_{x}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))\right)~.$ (3)
This strategy raises a privacy concern, as we exemplified in Section 3: Since $\mathbf{y}^{*}(\mathbf{x}),\mathbf{y}^{\lambda}(\mathbf{x})$ are required in order to compute $\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x})$, and are defined as the minimizers of $g(\mathbf{x},\cdot),\mathcal{L}_{\lambda}(\mathbf{x},\cdot)$ which are data-dependent, we cannot simply compute them under DP. In other words, even deciding where to invoke the gradient oracles, can already leak user information, hence breaking privacy before the gradients are even computed. We therefore must resort to approximating them using an auxiliary private method, for which we use DP-Loc(alized)-GD (Algorithm 2). [^5] We can replace the inner solver by any DP method that guarantees with high probability the optimal rate for strongly-convex objectives, as we further discuss in Appendix B.
In our analysis, we crucially rely on the fact that $g(\mathbf{x},\cdot)$ $\mathcal{L}_{\lambda}(\mathbf{x},\cdot)$ are both strongly-convex, which implies that optimizing them produces ${\widetilde{\mathbf{y}}}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t}$ such that the distances to the minimizers, namely $\|{\widetilde{\mathbf{y}}}_{t}-\mathbf{y}^{*}(\mathbf{x}_{t})\|,\|{\widetilde{\mathbf{y}}^{\lambda}_{t}-\mathbf{y}^{\lambda}(\mathbf{x}_{t})}\|$, are small. The distance bound is key, since then Eq. (3) allows using the smoothness of $f,g$ to translate the distance bounds into an inexact (i.e., biased) gradient oracle for $\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x}_{t})$, computed at the private points ${\widetilde{\mathbf{y}}}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t}$. We are able to bound the gradient bias induced by privatizing the inner solvers, as follows:
Lemma 4.5.
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}},\frac{L^{f}_{1}}{L^{g}_{1}}\}$, then the random variables $\widetilde{\mathbf{y}}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t}$ as defined in Algorithm 1 satisfy for all $t<T$ with probability at least $1-\gamma:$
$\left\|\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x}_{t})-\left[\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t})\right)\right]\right\|\leq\beta=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}\right)~.$
Having constructed an inexact hypergradient oracle, we can privatize its response using the standard Gaussian mechanism. To do so, we recall that the noise variance required to ensure privacy is tied to the component Lipschitz constants, and note that $\mathcal{L}_{\lambda}^{*}(\mathbf{x})$ decomposes as the finite-sum
$\mathcal{L}_{\lambda}^{*}(\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}^{*}_{\lambda,i}(\mathbf{x})~,~~~~\mathcal{L}^{*}_{\lambda,i}(\mathbf{x}):=f(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x});\xi_{i})+\lambda\left[g(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x});\xi_{i})-g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x});\xi_{i})\right]~.$
Bounding the Lipschitz constant of $\mathcal{L}_{\lambda,i}^{*}$ naively by applying the chain rule would result in the bound $\mathrm{Lip}(\mathbf{y}^{*})(L_{0}^{f}+2\lambda L_{0}^{g})\lesssim\lambda\mathrm{Lip}(\mathbf{y}^{*})L_{0}^{g}$, where $\mathrm{Lip}(\mathbf{y}^{*})$ is the Lipschitz constant of $\mathbf{y}^{*}(\mathbf{x}):\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{y}}$. Unfortunately, this bound grows with the penalty parameter $\lambda$, which will eventually be set large, and in particular, will grow with the dataset size $n$. We therefore need to derive a more nuanced analysis that allows obtaining a significantly tighter Lipschitz bound. We achieve this by showing that $\|\mathbf{y}^{\lambda}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\|\lesssim 1/\lambda$, which cancels out the multiplication by $\lambda$, and thus prove the following tight Lipschitz bound:
Lemma 4.6.
$\mathcal{L}^{*}_{\lambda,i}$ is $\mathcal{O}(\ell\kappa)$-Lipschitz independently of $\lambda$.
Finally, having constructed a private inexact stochastic oracle response for the smooth approximation $\mathcal{L}^{*}_{\lambda}$, we analyze an outer loop (update of $\mathbf{x}_{t}$), showing that is optimally robust to inexact and noisy gradients. Moreover, we employ an output rule which makes use of the already-privatized iterates, thus overcoming the need of additional noise in choosing the smallest gradient. This is a subtle technical challenge which is unique to the private setting, since non-private algorithms can simply validate the gradient norm and return the minimal one, yet in our case, using the Laplace mechanism to do so would spoil the convergence rate. In particular, we show that the corresponding trajectory and output rule lead a point with small (projected-)gradient norm, as stated below:
Proposition 4.7.
Suppose $h:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-smooth, that $\|\widetilde{\nabla}h(\cdot)-\nabla h(\cdot)\|\leq\beta$, and consider the following update rule with $\eta=\frac{1}{2L}:\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\big\{\big\langle\widetilde{\nabla}h(\mathbf{x}_{t})+\nu_{t},\mathbf{u}\big\rangle+\frac{1}{2\eta}\left\|\mathbf{x}_{t}-\mathbf{u}\right\|^{2}\big\}~,~\nu_{t}\sim\mathcal{N}(0,\sigma^{2}I)$ with the output rule $\mathbf{x}_{\mathrm{out}}:=\mathbf{x}_{t_{\mathrm{out}}},~t_{\mathrm{out}}:={\arg\min}_{t\in\{0,\dots,T-1\}}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|$. If $\alpha>0$ is such that $\alpha\geq C\max\{\beta,\sigma\sqrt{d\log(T/\gamma)}\}$ for a sufficiently large absolute constant $C>0$, then with probability at least $1-\gamma:~\left\|\mathcal{G}_{h,\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha$ for $T=\mathcal{O}\big(\frac{L(h(\mathbf{x}_{0})-\inf h)}{\alpha^{2}}\big)$.
Notably, for any $\alpha>0$, the result above allows the bias $\beta$ to be as large as $\Omega(\alpha)$ at all points, and the algorithm still reaches an $\alpha$-stationary point. Overall, applying Proposition 4.7 to $h=\mathcal{L}_{\lambda}^{*}$, we see that the (projected-)gradient norm can be as small as $\max\{\beta,\sigma\sqrt{d_{x}}\}$, up to logarithmic terms. Accounting for the smallest possible inexactness $\beta$ and noise addition $\sigma$ that ensure the the inner and outer loops, respectively, are both sufficiently private, we conclude the proof of Theorem 4.1; the full details appear in Section 8.
5 Mini-batch algorithm for DP bilevel ERM
In this section, we consider once again the bilevel ERM problem (BO/ERM), and provide Algorithm 3, which is a mini-batch variant of the previously discussed bilevel ERM algorithm. Given a mini-batch $B\subseteq S=\{\xi_{1},\dots,\xi_{n}\}$ and a function $h:\mathbb{R}^{d}\times\Xi\to\mathbb{R}$, we let $\nabla h(\mathbf{z};B)=\frac{1}{|B|}\sum_{\xi_{i}\in B}\nabla h(\mathbf{z};\xi_{i})$ denote the mini-batch gradient. We prove the following result:
Theorem 5.1.
Assume 2.1 and 2.2 hold, and that $\alpha\leq\ell\kappa^{3}\min\{\frac{1}{2\kappa},\frac{L_{0}^{g}}{L_{0}^{f}},\frac{L_{1}^{g}}{L_{1}^{f}},\frac{\Delta_{F}}{\ell\kappa}\}$. Then running Algorithm 3 with parameters assigned as in Theorem 4.1 and any batch sizes $b_{\mathrm{in}},b_{\mathrm{out}}\in[n]$ for sampling gradients, satisfies $(\epsilon,\delta)$-DP and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
$\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha=\widetilde{\mathcal{O}}\bigg(K_{1}\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+K_{2}\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}+K_{3}\cdot\frac{1}{b_{\mathrm{out}}}\bigg)~,$
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{3/4}\kappa^{5/4}),~K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{5/6}\kappa^{11/6}),~K_{3}=\mathcal{O}(\ell\kappa)$.
Remark 5.2 (Outer batch size dependence).
Algorithm 3 ensures privacy for any batch sizes, yet notably, the guaranteed gradient norm bound does not go to zero (as $n\to\infty$) for constant outer-batch size. The same phenomenon also holds for for “single”-level constrained nonconvex optimization, as noted by Ghadimi et al. (2016) (specifically, see Corollary 4 and related discussion therein). Accordingly, the inner-batch size $b_{\mathrm{in}}$ can be set whatsoever, while setting $b_{\mathrm{out}}=\mathcal{O}(\max\{(\epsilon n/\sqrt{d_{x}})^{1/2},(\epsilon n/\sqrt{d_{y}})^{1/3}\})\ll n$ recovers the full-batch rate (in general, $\|\mathcal{G}(\mathbf{x}_{\mathrm{out}})\|\overset{n\to\infty}{\longrightarrow}0$ whenever $b_{\mathrm{out}}\overset{n\to\infty}{\longrightarrow}\infty$). It is interesting to note that the $1/b_{\mathrm{out}}$ term shows up in the analysis only as an upper bound on the sub-Gaussian norm of the mini-batch gradient estimator. Thus, in applications for which some (possibly small) batch size results in reasonably accurate gradients, the result above holds with the outer mini-batch gradient’s standard deviation replacing $1/b_{\mathrm{out}}$, which is to be expected anyhow for high probability guarantees.
1:Input: Initialization $(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$, privacy budget $(\epsilon,\delta)$, penalty $\lambda>0$, noise level $\sigma^{2}>0$, step size $\eta>0$, iteration budget $T\in\mathbb{N}$, batch sizes $b_{\mathrm{in}},b_{\mathrm{out}}\in\mathbb{N}$.
2:for $t=0,\dots,T-1$ do
3:  Apply $(\frac{\epsilon}{\sqrt{18T}},\frac{\delta}{3(T+1)})$-DP-Loc-SGD (Algorithm 4) to solve $\triangleright$ Strongly-convex problems
${\widetilde{\mathbf{y}}}_{t}\approx\arg\min_{\mathbf{y}}g(\mathbf{x}_{t},\mathbf{y})~,~~~~~\widetilde{\mathbf{y}}^{\lambda}_{t}\approx\arg\min_{\mathbf{y}}\left[f(\mathbf{x}_{t},\mathbf{y})+\lambda\cdot g(\mathbf{x}_{t},\mathbf{y})\right]$
4:  $\widetilde{\mathbf{g}}_{t}=\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t};B_{t})\right)+\nu_{t},~B_{t}\sim S^{b_{\mathrm{out}}},~\nu_{t}\sim\mathcal{N}(\mathbf{0},\sigma^{2}I_{d_{x}})$
5:  $\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\left\langle\widetilde{\mathbf{g}}_{t},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}_{t}\right\|^{2}\right\}$ $\triangleright$ If $\mathcal{X}=\mathbb{R}^{d_{x}}$, then $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\widetilde{\mathbf{g}}_{t}$
6:end for
[…]
8:Output: $\mathbf{x}_{t_{\mathrm{out}}}$.
Algorithm 3 Mini-batch DP Bilevel
1:Input: Objective $h:\mathbb{R}^{d_{y}}\times\Xi\to\mathbb{R}$, initialization $\mathbf{y}_{0}\in\mathcal{Y}$, privacy budget $(\epsilon^{\prime},\delta^{\prime})$, batch size $b_{\mathrm{in}}\in\mathbb{N}$, number of rounds $M\in\mathbb{N}$, noise level $\sigma_{\mathrm{SGD}}^{2}>0$, step sizes $(\eta_{t})_{t=0}^{T-1}$, iteration budget $T_{\mathrm{SGD}}\in\mathbb{N}$, radii $(R_{m})_{m=0}^{M-1}>0$.
[…]
3:for $m=0,\dots,M-1$ do
4:  for $t=0,\dots,T_{\mathrm{SGD}}-1$ do
5:   $\mathbf{y}_{t+1}^{m}=\mathrm{Proj}_{\mathbb{B}(\mathbf{y}_{0}^{m},R_{m})}\left[\mathbf{y}_{t}^{m}-\eta_{t}\left(\nabla h(\mathbf{y}_{t}^{m};B_{t})+\nu_{t}\right)\right],~~B_{t}^{m}\sim S^{b_{\mathrm{in}}},~\nu_{t}^{m}\sim\mathcal{N}(\mathbf{0},\sigma_{\mathrm{SGD}}^{2}I_{d_{y}})$
6:  end for
7:  $\mathbf{y}_{0}^{m+1}=\frac{1}{T}\sum_{t=0}^{T-1}\mathbf{y}_{t}^{m}$
8:end for
9:Output: $\mathbf{y}_{\mathrm{out}}=\mathbf{y}^{M}_{0}$.
Algorithm 4 DP-Loc-SGD
The difference between Algorithm 3 and Algorithm 1, is that both the inner and outer loops sample mini-batch gradients. The inner loop guarantee is the same regardless of the inner-batch size $b_{\mathrm{in}}$, since for strongly-convex objectives it is possible to prove the same convergence rate guarantee for DP optimization in any case (as further discussed in Appendix B). As to the outer loop ($\mathbf{x}_{t+1}$ update rule), we apply standard concentration bounds to argue about the quality of the gradient estimates — hence the additive $1/b_{\mathrm{out}}$ factor — and rely on our analysis of the outer loop with inexact gradients (which are now even less exact due to sampling stochasticity). We remark that compared to the well-known analysis of Ghadimi et al. (2016) for mini-batch constrained nonconvex optimization, we derive high probability bounds without requiring several re-runs of the algorithm. We further remark that mini-batch sampling with replacement is analyzed for simplicity, though the same guarantees (up to constants) can be derived for sampling without replacement, at the cost of a more involved analysis.
6 Generalizing from ERM to population loss
We now move to consider stochastic (population) objectives, namely the problem (BO/Pop). We denote the population hyperobjective by $F_{\mathcal{P}}$, and as before $F_{S}$ denotes the empirical objective, where $S\sim\mathcal{P}^{n}$. We prove the following result:
Theorem 6.1.
Under Assumptions 2.1 and 2.2, if the preconditions of Theorem 4.1 hold, then Algorithm 1 is $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
$\left\|\mathcal{G}_{F_{\mathcal{P}},\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha=\widetilde{\mathcal{O}}\bigg(K_{1}\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+K_{2}\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}+\mathrm{gen}_{n}\bigg)~,$
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{1/4}\kappa^{5/4}),~K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{5/6}\kappa^{11/6})$, and
$\mathrm{gen}_{n}=\widetilde{\mathcal{O}}\bigg(\frac{\ell\kappa(\ell\kappa+1)(\sqrt{d_{x}}+\kappa)}{\sqrt{n}}+\frac{\ell^{1/2}\kappa^{3/2}(\ell\kappa+1)}{n^{1/4}}\bigg)~.$
Similarly, if the preconditions of Theorem 5.1 hold, then for any batch sizes $b_{\mathrm{in}},b_{\mathrm{out}}\in[n]$, Algorithm 3 is $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
$\left\|\mathcal{G}_{F_{\mathcal{P}},\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha=\widetilde{\mathcal{O}}\bigg(K_{1}\Big(\frac{\sqrt{d_{x}}}{\epsilon n}\Big)^{1/2}+K_{2}\Big(\frac{\sqrt{d_{y}}}{\epsilon n}\Big)^{1/3}+\frac{\ell\kappa}{b_{\mathrm{out}}}+\mathrm{gen}_{n}\bigg)~.$
The proof of Theorem 6.1 relies on establishing uniform convergence of the empirical hypergradients to the population hypergradients (Lemma 8.9), which as far as we know does not appear in prior literature and therefore may be of independent interest. This further implies uniform convergence of projected gradients, since the gradient mapping is non-expansive.
7 Application: private regularization hyperparameter tuning
In this section, we specialize our private bilevel framework to the well-studied problem of tuning a regularization hyperparameter. Specifying our analysis to this task results in a simple private algorithm, which as we will soon explain, offers an alternative paradigm for private hyperparameter tuning compared to existing works on this topic (Liu and Talwar, 2019; Papernot and Steinke, 2022).
We start by introducing the setting. Given a labeled dataset $S=(\mathbf{a}_{i},b_{i})_{i=1}^{n}\subset\mathbb{R}^{d_{a}}\times\mathbb{R}$, we aim to fit a parametric model $\{h_{\theta}:\mathbb{R}^{d_{a}}\to\mathbb{R}\mid\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}\}$ (e.g., linear model, or neural network) with respect to a loss function $\ell oss:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{\geq 0}$. A classic approach is to split the data into two disjoint datasets, train and validation $S=S_{\mathrm{tr}}\cup S_{\mathrm{val}}$, choose a regularizer $R:\Theta\to\mathbb{R}_{\geq 0}$ (e.g., $R(\theta)=\|\theta\|_{2}$ or $R(\theta)=\|\theta\|_{1}$) and minimize the validation loss of a model trained under regularization, while searching the best amount of regularization. This corresponds to the bilevel problem
$\mathrm{minimize}_{\omega\geq 0}~~~~F(\omega)$ $:=f(\omega,\theta^{*}_{\omega})=\frac{1}{|S_{\mathrm{val}}|}\sum_{(\mathbf{a}_{i},b_{i})\in S_{\mathrm{val}}}\ell oss(h_{\theta^{*}_{\omega}}(\mathbf{a}_{i}),b_{i})$
$\mathrm{subject~to}~~~~~~~~~\theta^{*}_{\omega}$ $:=\arg\min_{\theta\in\Theta}g(\omega,\theta)=\frac{1}{|S_{\mathrm{tr}}|}\sum_{(\mathbf{a}_{i},b_{i})\in S_{\mathrm{tr}}}\ell oss(h_{\theta}(\mathbf{a}_{i}),b_{i})+\omega\cdot R(\theta)~.$ (4)
The so-called model selection problem described above is extremely well studied, and strategies to solve it generally fall into two categories. The first approach is to discretize the domain of $\omega$ to a finite set $\Omega$, train models associated to each $\omega\in\Omega$, and pick the model minimizing the validation loss (cf. Hastie et al., 2009, Section 7 for a detailed discussion). This raises a privacy issue, since even if each model is private, the very choice of the hyperparameter can break privacy. Therefore, privatizing this approach requires employing a private selection algorithm over the hyperparameter, as studied by Liu and Talwar (2019); Papernot and Steinke (2022).
A second approach to hyperparameter tuning is based on differentiable programming: instead of predetermining a finite set of possible values for $\omega$, we can solve Eq. (4) using BO methods, namely “differentiate through $\omega$”, and update it accordingly on the fly. This approach was pioneered by Bengio (2000), and was further developed by multiple works (Maclaurin et al., 2015; Franceschi et al., 2017, 2018; Lorraine et al., 2020; Engstrom et al., 2025). However, we are not aware of a private variant of this methodology, and our goal is to provide one here.
Applying our algorithm to this problem turns out to carry some useful simplifications. Note that $\nabla_{\omega}f=0$, $\nabla_{\omega}g(\omega,\theta)=R(\theta)$, thus letting $\theta_{t}\approx\arg\min_{\mathbf{w}}g(\omega_{t},\theta)$ and $\theta_{t}^{\lambda}\approx\arg\min_{\theta}(f(\omega_{t},\theta)+\lambda g(\omega_{t},\theta))$ be private solutions to the inner problems, [^6] Note that $\theta_{t}$ can be interpreted as a regularized model over the training data, while $\theta_{t}^{\lambda}$ simply corresponds to a model trained by mixing in a bit of the validation loss (weighted down by $1/\lambda$). our algorithm yields the following outer loop:
$\omega_{t+1}$ $=\arg\min_{u\geq 0}\Big\{u\left[\nabla_{\omega}f(\omega_{t},\theta_{t}^{\lambda})+\lambda(\nabla_{\omega}g(\omega_{t},\theta_{t}^{\lambda})-\nabla_{\omega}g(\omega_{t},\theta_{t}))+\mathcal{N}(0,\sigma^{2})\right]+\tfrac{1}{2\eta}(u-\omega_{t})^{2}\Big\}$
$=\arg\min_{u\geq 0}\left\{u\left[0+\lambda(R(\theta_{t}^{\lambda})-R(\theta_{t}))+\mathcal{N}(0,\sigma^{2})\right]+\tfrac{1}{2\eta}(u-\omega_{t})^{2}\right\}$
$=\max\left\{0,~\omega_{t}+\eta\lambda\cdot\mathcal{N}\left(R(\theta_{t}^{\lambda})-R(\theta_{t}),\,\sigma^{2}/\lambda^{2}\right)\right\}~,$
where the last equality is easily verified by solving a one-dimensional quadratic over a half-line.
Overall, we obtain a simple closed-form update rule for privately tuning the regularization parameter $\omega$, based on two private training subroutines (which can be solved in a black-box manner), according to a normal variable centered at the difference between their associated complexities measured by $R$.
8 Proofs
Throughout the proof section, we abbreviate $f_{i}(\cdot)=f(\cdot\,;\xi_{i}),~g_{i}(\cdot)=g(\cdot\,;\xi_{i}),~F=F_{S}$.
8.1 Proof of Lemma 4.5
Note that the two sub-problems solved by Algorithm 1 are strongly-convex and admit Lipschitz components over $\mathcal{Y};~g(\mathbf{x},\cdot)$ by assumption, and $f+\lambda g$ by combining this with the smoothness/Lipschitzness of $f$, as follows:
Lemma 8.1.
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}}\}$ then for all $\mathbf{x}\in\mathcal{X}:~f(\mathbf{x},\cdot)+\lambda g(\mathbf{x},\cdot)$ is $\frac{\lambda\mu_{g}}{2}$ strongly-convex, and moreover for all $i\in[n]:~f_{i}(\mathbf{x},\cdot)+\lambda g_{i}(\mathbf{x},\cdot)$ is $2\lambda L_{0}^{g}$-Lipschitz.
We therefore invoke the following guarantee, which provides the optimal result for strongly-convex DP ERM via DP-Loc-GD (Algorithm 2).
Theorem 8.2.
Suppose that $h:\mathbb{R}^{d_{y}}\to\mathbb{R}$ is a $\mu$-strongly-convex function of the form $h(\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}h(\mathbf{y},\xi_{i})$ where $h(\cdot,\xi_{i})$ is $L$-Lipschitz for all $i\in[n]$. Suppose $\arg\min h=:\mathbf{y}^{*}\in\mathbb{B}(\mathbf{y}_{0},R_{0})$ and that $n\geq\frac{LR_{0}^{2/\log(d_{y})}}{\mu\epsilon^{\prime}}$. Then there is an assignment of parameters $M=\log_{2}\log(\frac{\mu\epsilon^{\prime}n}{L})\,,~\sigma^{2}_{\mathrm{GD}}=\widetilde{\mathcal{O}}(L^{2}/\epsilon^{\prime 2})\,,~\eta_{t}=\frac{1}{\mu(t+1)},~T_{\mathrm{GD}}=n^{2}\,,~R_{m}=\widetilde{\Theta}\left(\sqrt{\frac{R_{m-1}L}{\mu\epsilon^{\prime}n}}+\frac{L\sqrt{d_{y}}}{\mu\epsilon^{\prime}n}\right)$ such that running Algorithm 2 satisfies $(\epsilon^{\prime},\delta^{\prime})$-DP, and outputs $\mathbf{y}_{\mathrm{out}}$ such that $\left\|\mathbf{y}_{\mathrm{out}}-\mathbf{y}^{*}\right\|=\widetilde{\mathcal{O}}\left(\frac{L\sqrt{d_{y}}}{\mu n\epsilon^{\prime}}\right)$ with probability at least $1-\gamma$.
Although the rate in Theorem 8.2 appears in prior works such as (Bassily et al., 2014; Feldman et al., 2020), it is typically manifested through a bound in expectation (and in terms of function value) as opposed to with high probability, required for our purpose. We therefore, for the sake of completeness, provide a self-contained proof of Theorem 8.2 in Appendix B.
Applied to the functions $g(\mathbf{x}_{t},\cdot)$ and $f(\mathbf{x}_{t},\cdot)+\lambda g(\mathbf{x}_{t},\cdot)$, and invoking Lemma 8.1, yields the following.
Corollary 8.3.
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}}\}$, then $\widetilde{\mathbf{y}}_{t}$ and $\widetilde{\mathbf{y}}_{t}^{\lambda}$ (as appear in Algorithm 1) satisfy with probability at least $1-\gamma:$
$\max\left\{\left\|\widetilde{\mathbf{y}}_{t}-\mathbf{y}^{*}(\mathbf{x}_{t})\right\|,~\|\widetilde{\mathbf{y}}_{t}^{\lambda}-\mathbf{y}^{\lambda}(\mathbf{x}_{t})\|\right\}=\widetilde{\mathcal{O}}\left(\frac{L_{0}^{g}\sqrt{d_{y}T}}{\epsilon\mu_{g}n}\right)~.$
We are now ready to prove the main proposition of this section, which we restate below:
See 4.5
Proof of Lemma 4.5.
As in Eq. (3), we note that by construction $\mathcal{L}_{\lambda}^{*}(\mathbf{x})=\arg\min_{\mathbf{y}}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y})$, therefore it holds that
$\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x})=\nabla_{x}f(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))+\lambda\left(\nabla_{x}g(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))-\nabla_{x}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))\right)~.$
Denoting $\mathbf{g}_{t}=\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t})\right)$, by the smoothness of $f$ and $g$ we see that
$\left\|\mathbf{g}_{t}-\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x}_{t})\right\|\leq L_{1}^{f}\left\|\widetilde{\mathbf{y}}_{t}^{\lambda}-\mathbf{y}^{\lambda}(\mathbf{x}_{t})\right\|+\lambda L_{1}^{g}\left\|\widetilde{\mathbf{y}}_{t}^{\lambda}-\mathbf{y}^{\lambda}(\mathbf{x}_{t})\right\|+\lambda L_{1}^{g}\left\|\widetilde{\mathbf{y}}_{t}-\mathbf{y}^{*}(\mathbf{x}_{t})\right\|~.$
Applying Corollary 8.3 and union bounding over $T$, we can further bound the above as
$\left\|\mathbf{g}_{t}-\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x}_{t})\right\|$ $=\widetilde{\mathcal{O}}\left(\frac{L_{1}^{f}L_{0}^{g}\sqrt{d_{y}T}}{\epsilon\mu_{g}n}+\frac{\lambda L_{1}^{g}L_{0}^{g}\sqrt{d_{y}T}}{\epsilon\mu_{g}n}+\frac{\lambda L_{1}^{g}L_{0}^{g}\sqrt{d_{y}T}}{\epsilon\mu_{g}n}\right)$
$=\widetilde{\mathcal{O}}\left(\frac{\lambda L_{1}^{g}L_{0}^{g}\sqrt{d_{y}T}}{\epsilon\mu_{g}n}\right)$
$=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}\right)~,$
where the second bound holds for $\lambda\geq\frac{L_{1}^{f}}{L_{1}^{g}}$ . ∎
8.2 Proof of Lemma 4.6
We start by providing two lemmas, both of which borrow ideas that appeared in the smoothness analysis of Chen et al. (2024).
Lemma 8.4.
$\mathbf{y}^{\lambda}(\mathbf{x}):\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{y}}$ is $\left(\frac{4L_{1}^{g}}{\mu_{g}}\right)$-Lipschitz.
Proof of Lemma 8.4.
Differentiating $\nabla_{y}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))=\mathbf{0}$ with respect the first argument gives
$\nabla^{2}_{xy}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))+\nabla\mathbf{y}^{\lambda}(\mathbf{x})\cdot\nabla_{yy}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))=\mathbf{0}~,$
hence
$\nabla\mathbf{y}^{\lambda}(\mathbf{x})=-\nabla^{2}_{xy}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))\cdot\left[\nabla_{yy}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))\right]^{-1}~.$
Noting that $\nabla^{2}_{xy}\mathcal{L}_{\lambda}\preceq 2\lambda L_{1}^{g}$ and $\nabla^{2}_{yy}\mathcal{L}_{\lambda}\succeq\lambda\mu_{g}/2$ everywhere, hence $\nabla^{2}_{yy}\mathcal{L}_{\lambda}]^{-1}\preceq 2/\lambda\mu_{g}$ we get that
$\left\|\nabla\mathbf{y}^{\lambda}(\mathbf{x})\right\|\leq 2\lambda L_{1}^{g}\cdot\frac{2}{\lambda\mu_{g}}=\frac{4L_{1}^{g}}{\mu_{g}}~.$
Lemma 8.5.
For all $\mathbf{x}\in\mathcal{X}:~\left\|\mathbf{y}^{\lambda}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\right\|\leq\frac{L_{0}^{f}}{\lambda\mu_{g}}$.
Proof of Lemma 8.5.
First, note that by definition of $\mathbf{y}^{\lambda}(\mathbf{x})$ it holds that
$\mathbf{0}=\nabla_{y}\mathcal{L}_{\lambda}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))=\nabla_{y}f(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))+\lambda\nabla_{y}g(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))~,$
hence
$\nabla_{y}g(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))=-\frac{1}{\lambda}\cdot\nabla_{y}f(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))~,$
so in particular by the Lipschitz assumption on $f$ we see that
$\left\|\nabla_{y}g(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))\right\|\leq\frac{L_{0}^{f}}{\lambda}~.$
By invoking the $\mu$-strong convexity of $g$ we further get
$\left\|\mathbf{y}^{\lambda}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\right\|\leq$~ $\frac{1}{\mu}\left\|\nabla_{y}g(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))-\underset{=\mathbf{0}}{\underbrace{\nabla_{y}g(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))}}\right\|\leq\frac{L_{0}^{f}}{\lambda\mu}~.$
We are now ready to prove the main result of this section, restated below.
See 4.6
Proof of Lemma 4.6.
For all $\mathbf{x}\in\mathcal{X}$ it holds that
$\left\|\nabla\mathcal{L}^{*}_{\lambda,i}(\mathbf{x})\right\|$ $=\left\|\nabla_{x}f_{i}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))+\lambda\left[\nabla_{x}g_{i}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))-\nabla_{x}g_{i}(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))\right]\right\|$
$\leq\left\|\nabla_{x}f_{i}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))\right\|+\lambda\left\|\nabla_{x}g_{i}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))-\nabla_{x}g_{i}(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))\right\|~,$ (5)
thus we will bound each of the summands above.
For the first term, since $\mathbf{y}^{\lambda}$ is $\frac{4L_{1}^{g}}{\mu_{g}}$ Lipschitz according to Lemma 8.4, and $f_{i}$ is $L_{0}^{f}$-Lipschitz by assumption, the chain rule yields the bound
$\left\|\nabla_{x}f_{i}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))\right\|\leq\frac{4L_{1}^{g}L_{0}^{f}}{\mu_{g}}\leq 4\ell\kappa~.$ (6)
As to the second term, since $g_{i}$ is $L_{1}^{g}$-smooth, we use Lemma 8.5 and get that
$\lambda\left\|\nabla_{x}g_{i}(\mathbf{x},\mathbf{y}^{\lambda}(\mathbf{x}))-\nabla_{x}g_{i}(\mathbf{x},\mathbf{y}^{*}(\mathbf{x}))\right\|\leq\lambda L_{1}^{g}\left\|\mathbf{y}^{\lambda}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\right\|\leq\frac{L_{1}^{g}L_{0}^{f}}{\mu_{g}}\leq\ell\kappa~.$ (7)
Plugging Eqs. (6) and (7) into Eq. (5) completes the proof.
8.3 Proof of Proposition 4.7
As $\nu_{0},\dots,\nu_{T-1}\overset{iid}{\sim}\mathcal{N}(0,\sigma^{2}I)$, a standard Gaussian norm bound (cf. Vershynin 2018, Theorem 3.1.1) ensures that with probability at least $1-\gamma$, for all $t\in\{0,1,\dots,T-1\}:~\left\|\nu_{t}\right\|^{2}\lesssim d\sigma^{2}\log(T/\gamma)\lesssim\frac{\alpha^{2}}{64}$. We therefore condition the rest of the proof on the highly probable event under which this uniform norm bound indeed holds. We continue by introducing some notation. We denote $\widetilde{\nabla}_{t}=\widetilde{\nabla}{h}(\mathbf{x}_{t})+\nu_{t}$, and $\delta_{t}:=\widetilde{\nabla}_{t}-\nabla h(\mathbf{x}_{t})$. We further denote
$\mathbf{x}_{t}^{+}$ $:=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\left\langle\nabla h(\mathbf{x}_{t}),\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{x}_{t}-\mathbf{u}\right\|^{2}\right\}~,$
$\mathcal{G}_{t}$ $:=\frac{1}{\eta}(\mathbf{x}_{t}-\mathbf{x}_{t}^{+})~,$
$\rho_{t}$ $:=\frac{1}{\eta}(\mathbf{x}_{t}-\mathbf{x}_{t+1})~.$
Note that by construction,
$\mathcal{G}_{t}=\mathcal{G}_{h,\eta}(\mathbf{x}_{t}):=\frac{1}{\eta}\left(\mathbf{x}_{t}-\mathcal{P}_{\nabla h,\eta}(\mathbf{x}_{t})\right)~,~~~~~\mathcal{P}_{\nabla h,\eta}(\mathbf{x}_{t}):=\arg\min_{\mathbf{u}\in\mathcal{X}}\left[\left\langle\nabla h(\mathbf{x}_{t}),\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}_{t}\right\|^{2}\right]~,$
and that $\mathcal{G}_{t_{\mathrm{out}}}$ is precisely the quantity we aim to bound. We start by proving some useful lemmas regarding the quantities defined above.
Lemma 8.6.
Under the event that $\left\|\nu_{t}\right\|^{2}\leq\frac{\alpha^{2}}{64}$ for all $t$, it holds that $\left\|\delta_{t}\right\|\leq\frac{\alpha}{4}$.
Proof.
By our assumptions on $\beta,\left\|\nu_{t}\right\|$, we get that
$\left\|\delta_{t}\right\|\leq\|\widetilde{\nabla}h(\mathbf{x}_{t})-\nabla h(\mathbf{x}_{t})\|+\left\|\nu_{t}\right\|\leq\beta+\frac{\alpha}{8}\leq\frac{\alpha}{4}~.$
Lemma 8.7.
It holds that $\langle\widetilde{\nabla}_{t},\rho_{t}\rangle\geq\left\|\rho_{t}\right\|^{2}$.
Proof.
By definition, $\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\langle\widetilde{\nabla}_{t},\mathbf{u}\rangle+\frac{1}{2\eta}\left\|\mathbf{x}_{t}-\mathbf{u}\right\|^{2}\right\}$. Hence, by the first-order optimality criterion, for any $\mathbf{u}\in\mathcal{X}:$
$\left\langle\widetilde{\nabla}_{t}+\frac{1}{\eta}(\mathbf{x}_{t+1}-\mathbf{x}_{t}),\mathbf{u}-\mathbf{x}_{t+1}\right\rangle\geq 0~.$
In particular, setting $\mathbf{u}=\mathbf{x}_{t}$ yields
$0\leq\left\langle\widetilde{\nabla}_{t}+\frac{1}{\eta}(\mathbf{x}_{t+1}-\mathbf{x}_{t}),\mathbf{x}_{t}-\mathbf{x}_{t+1}\right\rangle=\left\langle\widetilde{\nabla}_{t}-\rho_{t},\eta\rho_{t}\right\rangle=\eta\left(\left\langle\widetilde{\nabla}_{t},\rho_{t}\right\rangle-\left\|\rho_{t}\right\|^{2}\right)~,$
which proves the claim since $\eta>0$. ∎
With the lemmas above in hand, we are now ready to prove Proposition 4.7. Note that by construction, the algorithm returns the index $t$ with minimal $\left\|\rho_{t}\right\|$. Further note that $\left\|\rho_{t}-\mathcal{G}_{t}\right\|\leq\left\|\delta_{t}\right\|$ by Lemma C.1, thus
$\left\|\mathcal{G}_{t}\right\|\leq\left\|\rho_{t}\right\|+\left\|\delta_{t}\right\|\leq\left\|\rho_{t}\right\|+\frac{\alpha}{4}~,$ (8)
where the last inequality is due to Lemma 8.6, hence it suffices to bound $\left\|\rho_{t_{\mathrm{out}}}\right\|$ (which is the quantity measured by the output rule). To that end, by smoothness, we have for any $t\in\{0,1,\dots,T-2\}:$
$h(\mathbf{x}_{t+1})$ $\leq h(\mathbf{x}_{t})+\left\langle\nabla h(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L}{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}$
$=h(\mathbf{x}_{t})-\eta\left\langle\nabla h(\mathbf{x}_{t}),\rho_{t}\right\rangle+\frac{L\eta^{2}}{2}\left\|\rho_{t}\right\|^{2}$
$=h(\mathbf{x}_{t})-\eta\left\langle\widetilde{\nabla}_{t},\rho_{t}\right\rangle+\frac{L\eta^{2}}{2}\left\|\rho_{t}\right\|^{2}+\eta\left\langle\delta_{t},\rho_{t}\right\rangle$
$\leq h(\mathbf{x}_{t})-\eta\left\|\rho_{t}\right\|^{2}+\frac{L\eta^{2}}{2}\left\|\rho_{t}\right\|^{2}+\eta\left\|\delta_{t}\right\|\cdot\left\|\rho_{t}\right\|~,$
where the last inequality followed by applying Lemma 8.7 and Cauchy-Schwarz. Rearranging, and recalling that $\eta=\frac{1}{2L}$, hence $1<2-L\eta$ and also $\frac{1}{\eta}=2L$, we get that
$\left\|\rho_{t}\right\|^{2}-2\left\|\delta_{t}\right\|\cdot\left\|\rho_{t}\right\|\leq\left(2-L\eta\right)\left\|\rho_{t}\right\|^{2}-2\left\|\delta_{t}\right\|\cdot\left\|\rho_{t}\right\|\leq\frac{2\left(h(\mathbf{x}_{t})-h(\mathbf{x}_{t+1})\right)}{\eta}=4L\left(h(\mathbf{x}_{t})-h(\mathbf{x}_{t+1})\right)~.$
Summing over $t\in\{0,1\dots,T-1\}$, using the telescoping property of the right hand side, and dividing by $T$ gives that
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|\rho_{t}\right\|\left(\left\|\rho_{t}\right\|-2\left\|\delta_{t}\right\|\right)\leq\frac{4L\left(h(\mathbf{x}_{0})-\inf h\right)}{T}~.$ (9)
Note that if for some $t\in\{0,1,\dots,T-1\}:~\left\|\rho_{t}\right\|\leq\frac{3\alpha}{4}$ then we have proved our desired claim by Eq. (8) and the fact that $\left\|\rho_{t_{\mathrm{out}}}\right\|=\min_{t}\left\|\rho_{t}\right\|$ by definition. On the other hand, assuming that $\left\|\rho_{t}\right\|>\frac{3\alpha}{4}$ for all $t$, invoking Lemma 8.6, we see that $\left\|\rho_{t}\right\|-2\left\|\delta_{t}\right\|\geq\left\|\rho_{t}\right\|-\frac{\alpha}{2}\geq\frac{1}{3}\left\|\rho_{t}\right\|$, which implies $\left\|\rho_{t}\right\|\left(\left\|\rho_{t}\right\|-2\left\|\delta_{t}\right\|\right)\geq\frac{1}{3}\left\|\rho_{t}\right\|^{2}$. Combining this with Eq. (9) yields
$\left\|\rho_{t_{\mathrm{out}}}\right\|^{2}=\min_{t\in\{0,1,\dots,T-1\}}\left\|\rho_{t}\right\|^{2}\leq\frac{1}{T}\sum_{t=0}^{T-1}\left\|\rho_{t}\right\|^{2}\leq\frac{12L\left(h(\mathbf{x}_{0})-\inf h\right)}{T}~,$
and the right side is bounded by $\frac{9\alpha^{2}}{16}$ for $T=\mathcal{O}\left(\frac{L(h(\mathbf{x}_{0})-\inf h)}{\alpha^{2}}\right)$, finishing the proof by Eq. (8).
8.4 Proof of Theorem 4.1
We start by proving the privacy guarantee. Since $\mathcal{L}^{*}_{\lambda,i}$ is $\mathcal{O}(\ell\kappa)$-Lipschitz by Lemma 4.6, the sensitivity of $\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t})\right)$ is at most $\mathcal{O}(\ell\kappa)$. Hence, by setting $\sigma^{2}=C\frac{\ell^{2}\kappa^{2}\log(T/\delta)T}{\epsilon^{2}n^{2}}$ for a sufficiently large absolute constant $C>0$, $\widetilde{\mathbf{g}}_{t}$ is $(\frac{\epsilon}{\sqrt{18T}},\frac{\delta}{3(T+1)})$-DP. By basic composition, since each iteration also runs $(\frac{\epsilon}{\sqrt{18T}},\frac{\delta}{3(T+1)})$-DP-Loc-GD twice, we see that each iteration of the algorithm is $(3\cdot\frac{\epsilon}{\sqrt{18T}},3\cdot\frac{\delta}{3(T+1)})=(\frac{\epsilon}{\sqrt{2T}},\frac{\delta}{(T+1)})$-DP. Noting that under our parameter assignment $\frac{\epsilon}{\sqrt{T}}\ll 1$, by advanced composition we get that throughout $T$ iterations, the algorithm is overall $(\epsilon,\delta)$-DP as claimed.
We turn to analyze the utility of the algorithm. It holds that
$\left\|\mathcal{G}_{F,\eta}(\mathbf{x}_{t_{\mathrm{out}}})\right\|$ $\leq\left\|\mathcal{G}_{F,\eta}(\mathbf{x}_{t_{\mathrm{out}}})-\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x})\right\|+\left\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\right\|$
$\leq\|\nabla F(\mathbf{x}_{t_{\mathrm{out}}})-\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x})\|+\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$
$\lesssim\frac{\ell\kappa^{3}}{\lambda}+\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$
$\leq\frac{\alpha}{2}+\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|~,$ (10)
where the second inequality is due to Lemma C.1, the third due to Lemma 4.4.b, and the last by our assignment of $\lambda$. It therefore remains to bound $\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$.
To that end, applying Proposition 4.7 to the function $h=\mathcal{L}^{*}_{\lambda}$, under our assignment of $T$ — which accounts for the smoothness and initial sub-optimality bounds ensured by Lemma 4.4 — we get that $\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|\leq\frac{\alpha}{2}$, for $\alpha$ as small as
$\alpha=\Theta\left(\max\{\beta,\sigma\sqrt{d_{x}\log(T/\gamma)}\}\right)$ (11)
By Lemma 4.5, it holds that
$\beta=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell^{5/2}\kappa^{11/2}\Delta_{F}^{1/2}\sqrt{d_{y}}}{\alpha^{2}\epsilon n}\right)$ (12)
and we also have
$\sigma\sqrt{d_{x}\log(T/\gamma)}=\widetilde{\mathcal{O}}\left(\frac{\ell^{3/2}\kappa^{5/2}\Delta_{F}^{1/2}\sqrt{d_{x}}}{\alpha\epsilon n}\right)~.$ (13)
Plugging (12) and (13) back into Eq. (11), and solving for $\alpha$, completes the proof.
8.5 Proof of Theorem 5.1
Throughout this section, we abbreviate $b=b_{\mathrm{out}}$. We will need the following lemma, which is the mini-batch analogue of Lemma 4.5 from the full-batch setting.
Lemma 8.8.
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}},\frac{L^{f}_{1}}{L^{g}_{1}}\}$, then there is $\beta_{b}=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}+\frac{\ell\kappa}{b}\right)$ such that with probability at least $1-\gamma/2$, $\mathbf{g}_{t}^{B}:=\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};B_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t};B_{t})\right)$ satisfies for all $t\in\{0,\dots,T-1\}:~\|\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x}_{t})-\mathbf{g}^{B}_{t}\|\leq\beta_{b}$.
Proof of Lemma 8.8.
It holds that
$\left\|\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x}_{t})-\mathbf{g}^{B}_{t}\right\|\leq\left\|\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x}_{t})-\mathbb{E}[\mathbf{g}^{B}_{t}]\right\|+\left\|\mathbf{g}^{B}_{t}-\mathbb{E}[\mathbf{g}^{B}_{t}]\right\|~.$
To bound the first summand, note that $\mathbb{E}[\mathbf{g}^{B}_{t}]=\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})+\lambda\left(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t})-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t})\right)$, and therefore with probability at least $1-\gamma/4:$
$\left\|\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x}_{t})-\mathbb{E}[\mathbf{g}^{B}_{t}]\right\|=\mathcal{O}\left(\frac{\lambda L_{1}^{g}L_{0}^{g}\sqrt{d_{y}T}}{\epsilon\mu_{g}n}\right)=\mathcal{O}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}\right)~,$
following the same proof as Lemma 4.5 in Section 8.1, by replacing Theorem 8.2 by the mini-batch Theorem B.1 (whose guarantee holds regardless of the inner batch size) .
To bound the second summand, note that $\|\nabla_{x}f(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};\xi)+\lambda(\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t};\xi)-\nabla_{x}g(\mathbf{x}_{t},\widetilde{\mathbf{y}}_{t};\xi))\|\leq M=\mathcal{O}(\ell\kappa)$ for every $\xi\in\Xi$, by Lemma 4.6. Hence, $\mathbf{g}^{B}_{t}$ is the average of $b$ independent vectors bounded by $M$, all with the same mean, and therefore a standard concentration bound (cf. Jin et al. 2019) ensures that $\|\mathbf{g}_{t}^{B}-\mathbb{E}[\mathbf{g}_{t}^{B}]\|=\widetilde{\mathcal{O}}(M/b)$ with probability at least $1-\gamma/4$, which completes the proof. ∎
We can now prove the main mini-batch result:
Proof of Theorem 5.1.
[…]
We turn to analyze the utility of the algorithm. It holds that
$\left\|\mathcal{G}_{F,\eta}(\mathbf{x}_{t_{\mathrm{out}}})\right\|$ $\leq\left\|\mathcal{G}_{F,\eta}(\mathbf{x}_{t_{\mathrm{out}}})-\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x})\right\|+\left\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\right\|$
$\leq\|\nabla F(\mathbf{x}_{t_{\mathrm{out}}})-\nabla\mathcal{L}^{*}_{\lambda}(\mathbf{x})\|+\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$
$\lesssim\frac{\ell\kappa^{3}}{\lambda}+\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$
$\leq\frac{\alpha}{2}+\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|~,$ (14)
where the second inequality is due to Lemma C.1, the third due to Lemma 4.4, and the last by our assignment of $\lambda$. It therefore remains to bound $\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|$.
To that end, applying Proposition 4.7 to the function $h=\mathcal{L}^{*}_{\lambda}$, under our assignment of $T$ — which accounts for the smoothness and initial sub-optimality bounds ensured by Lemma 4.4 — we get that $\|\mathcal{G}_{\mathcal{L}^{*}_{\lambda},\eta}(\mathbf{x}_{t_{\mathrm{out}}})\|\leq\frac{\alpha}{2}$, for $\alpha$ as small as
$\alpha=\Theta\left(\max\{\beta_{b},\sigma\sqrt{d_{x}\log(T/\gamma)}\}\right)$ (15)
By Lemma 8.8, it holds that
$\beta_{b}=\widetilde{\mathcal{O}}\left(\frac{\lambda\ell\kappa\sqrt{d_{y}T}}{\epsilon n}+\frac{\ell\kappa}{b}\right)=\widetilde{\mathcal{O}}\left(\frac{\ell^{5/2}\kappa^{11/2}\Delta_{F}^{1/2}\sqrt{d_{y}}}{\alpha^{2}\epsilon n}+\frac{\ell\kappa}{b}\right)~,$ (16)
and we also have
$\sigma\sqrt{d_{x}\log(T/\gamma)}=\widetilde{\mathcal{O}}\left(\frac{\ell^{3/2}\kappa^{5/2}\Delta_{F}^{1/2}\sqrt{d_{x}}}{\alpha\epsilon n}\right)~.$ (17)
Plugging (16) and (17) back into Eq. (15), and solving for $\alpha$, completes the proof.
8.6 Proof of Theorem 6.1
The proof is based on the following lemma that establishes uniform convergence of hypergradients.
Lemma 8.9.
Suppose $\overline{\mathcal{X}}\subset\mathbb{R}^{d}_{x}$ is a subset of bounded diameter $\mathrm{diam}(\overline{\mathcal{X}})\leq D$, and that $S\sim\mathcal{P}^{n}$. Then with probability at least $1-\gamma$ for all $\mathbf{x}\in\overline{\mathcal{X}}:$
[…]
$+\frac{(\ell^{3/2}\kappa^{5/2}+\ell^{1/2}\kappa^{3/2})\log^{1/4}(1/\gamma)}{n^{1/4}}\bigg)~.$
Proof of Lemma 8.9.
We start by setting up notation. Given a dataset $S\sim\mathcal{P}^{n}$, recall that we denote by $F_{\mathcal{P}}/F_{S}$ the stochastic/empirical hyperobjectives, by $f_{\mathcal{P}}/f_{S}$ and $g_{\mathcal{P}}/g_{S}$ the stochastic/empirical upper- and lower-level objectives. We denote by $\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})=\arg\min_{\mathbf{y}}g_{\mathcal{P}}(\mathbf{x},\mathbf{y})$ and $\mathbf{y}^{*}_{S}(\mathbf{x})=\arg\min_{\mathbf{y}}g_{S}(\mathbf{x},\mathbf{y})$. It holds that
$\nabla F_{\mathcal{P}}(\mathbf{x})$ $=\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))+(\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}})^{\top}\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))$
$\nabla F_{S}(\mathbf{x})$ $=\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))+(\frac{d\mathbf{y}^{*}_{S}(\mathbf{x})}{d\mathbf{x}})^{\top}\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))~.$
Thus,
$\|\nabla F_{\mathcal{P}}(\mathbf{x})-\nabla F_{S}(\mathbf{x})\|$ $\leq\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|$
$~~~+\|(\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}})^{\top}\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-(\frac{d\mathbf{y}^{*}_{S}(\mathbf{x})}{d\mathbf{x}})^{\top}\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|$
$\leq\underset{(1)}{\underbrace{\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|}}$
$~~~+\underset{(2)}{\underbrace{\|\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}}\|\cdot\|\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|}}$
$~~~+\underset{(3)}{\underbrace{\|\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|\cdot\|\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}}-\frac{d\mathbf{y}^{*}_{S}(\mathbf{x})}{d\mathbf{x}}\|}}~$,
where the last inequality used the fact that $|ab-cd|=|ab-ad+ad-cd|\leq|a||b-d|+|d||a-c|$. We turn to bound each summand with high probability, which immediately results in Lemma 8.9 by summing and union bounding.
Bound $(1)$.
By smoothness, it holds that
$\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|\leq\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))\|+L_{1}^{f}\|\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})-\mathbf{y}^{*}_{S}(\mathbf{x})\|~.$
To bound the first term, we note that the partial derivatives are bounded by $L_{0}^{f}$ by Lipschitzness. Hence, we can apply a uniform convergence bound due to Mei et al. (2018, Theorem 1) and get with probability at least $1-\gamma/9:$
$\|\nabla_{x}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{x}f_{S}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))\|=\widetilde{\mathcal{O}}\left(L_{0}^{f}\sqrt{{d_{x}\log(D/\gamma)}/{n}}\right)~.$
To bound the second term, we use strong-convexity to get
$L_{1}^{f}\|\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})-\mathbf{y}^{*}_{S}(\mathbf{x})\|\leq\frac{L_{1}^{f}}{\sqrt{\mu_{g}}}\sqrt{g_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-g_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))}~.$
Now, we note that since $\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}),\mathbf{y}^{*}_{S}(\mathbf{x})$ are the stochastic/empirical minima of the strongly-convex stochastic optimization problem given by $g_{\mathcal{P}}$, a result due to Feldman and Vondrak (2019) bounds the generalization gap with probability at least $1-\gamma/9$ by
$g_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-g_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))=\widetilde{\mathcal{O}}\left(\frac{(L_{0}^{g})^{2}\log(1/\gamma)}{\mu_{g}n}+\frac{\sqrt{\log(1/\gamma)}}{\sqrt{n}}\right)~.$
Overall, by union bounding and simplifying the condition dependent constants, we get
$(1)=\widetilde{\mathcal{O}}\left(\frac{\ell\sqrt{{d_{x}\log(D/\gamma)}}}{\sqrt{n}}+\frac{\ell\kappa\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{\sqrt{\ell\kappa}\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
Bound $(2)$.
On one hand, by the implicit function theorem,
$\|\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}}\|=\|\nabla^{2}_{xy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))[\nabla^{2}_{yy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))]^{-1}\|\leq L_{1}^{g}/\mu_{g}\leq\kappa~.$
Moreover, repeating the argument used to bound $(1)$, we get
$\|\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|$ $\leq\|\nabla_{y}f_{\mathcal{P}}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))-\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x}))\|+L_{1}^{f}\|\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})-\mathbf{y}^{*}_{S}(\mathbf{x})\|$
$=\widetilde{\mathcal{O}}\left(\frac{\ell\sqrt{{d_{x}\log(D/\gamma)}}}{\sqrt{n}}+\frac{\ell\kappa\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{\sqrt{\ell\kappa}\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
Overall, we get
$(2)=\widetilde{\mathcal{O}}\left(\frac{\ell\kappa\sqrt{{d_{x}\log(D/\gamma)}}}{\sqrt{n}}+\frac{\ell\kappa^{2}\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{\ell^{1/2}{\kappa}^{3/2}\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
Bound $(3)$.
We first note that $\|\nabla_{y}f_{S}(\mathbf{x},\mathbf{y}^{*}_{S}(\mathbf{x}))\|\leq L_{0}^{f}\leq\ell$. Moreover, due to the implicit function theorem, the basic estimate $|ab-cd|\leq|a||b-d|+|d||a-c|$, the fact $\|A^{-1}-B^{-1}\|\leq\frac{1}{\lambda_{\min}(A)\lambda_{\min}(B)}\|A-B\|$, and second-order smoothness of $g$, we get
$\|\frac{d\mathbf{y}^{*}_{\mathcal{P}}(\mathbf{x})}{d\mathbf{x}}-\frac{d\mathbf{y}^{*}_{S}(\mathbf{x})}{d\mathbf{x}}\|$ $=\|\nabla^{2}_{xy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))[\nabla^{2}_{yy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))]^{-1}-\nabla^{2}_{xy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))[\nabla^{2}_{yy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))]^{-1}\|$
$\leq\|\nabla^{2}_{xy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))\|\cdot\|[\nabla^{2}_{yy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))]^{-1}-[\nabla^{2}_{yy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))]^{-1}\|$
$~~~~+\|[\nabla^{2}_{yy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))]^{-1}\|\cdot\|\nabla^{2}_{xy}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))-\nabla^{2}_{xy}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))\|$
$\leq(\frac{L_{1}^{g}}{\mu_{g}^{2}}+\frac{1}{\mu_{g}})\|\nabla^{2}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))-\nabla^{2}g_{S}(\mathbf{x},\mathbf{y}_{S}^{*}(\mathbf{x}))\|$
$\leq(\frac{L_{1}^{g}}{\mu_{g}^{2}}+\frac{1}{\mu_{g}})(\|\nabla^{2}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))-\nabla^{2}g_{S}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))\|+L_{2}^{g}\|\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x})-\mathbf{y}_{S}^{*}(\mathbf{x})\|)~.$
Applying a Hessian uniform convergence bound due to Mei et al. (2018, Theorem 1.b) gives with probability at least $1-\gamma/9:$
$\|\nabla^{2}g_{\mathcal{P}}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))-\nabla^{2}g_{S}(\mathbf{x},\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x}))\|=\widetilde{\mathcal{O}}\left(L_{1}^{g}\sqrt{{d_{x}\log(D/\gamma)}/{n}}\right)~.$
The same argument that appeared in bounding $(1)$ and $(2)$ gives
$L_{2}^{g}\|\mathbf{y}_{\mathcal{P}}^{*}(\mathbf{x})-\mathbf{y}_{S}^{*}(\mathbf{x})\|=\widetilde{\mathcal{O}}\left(\frac{\ell\kappa\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{\sqrt{\ell\kappa}\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
Combining all the pieces and simplifying the condition dependent constants, we get overall
$(3)=\widetilde{\mathcal{O}}\left(\frac{(\ell^{2}\kappa^{2}+\ell\kappa)\sqrt{d_{x}\log(D/\gamma)}}{\sqrt{n}}+\frac{(\ell^{2}\kappa^{3}+\ell\kappa^{2})\sqrt{\log(1/\gamma)}}{\sqrt{n}}+\frac{(\ell^{3/2}\kappa^{5/2}+\ell^{1/2}\kappa^{3/2})\log^{1/4}(1/\gamma)}{n^{1/4}}\right)~.$
With Lemma 8.9 in hand, we turn to prove Theorem 6.1.
Proof of Theorem 6.1.
Using the fact that the gradient mapping is non-expansive (Lemma C.1), we see that with probability at least $1-\gamma/2$ the following holds:
$\left\|\mathcal{G}_{F_{\mathcal{P}},\eta}(\mathbf{x}_{\mathrm{out}})\right\|$ $\leq\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|+\left\|\mathcal{G}_{F_{\mathcal{P}},\eta}(\mathbf{x}_{\mathrm{out}})-\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|$
$\leq\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|+\left\|\nabla F_{\mathcal{P}}(\mathbf{x}_{\mathrm{out}})-\nabla F_{S}(\mathbf{x}_{\mathrm{out}})\right\|$
$=\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|+\widetilde{\mathcal{O}}\bigg(\frac{(\ell^{2}\kappa^{2}+\ell\kappa)\sqrt{d_{x}\log(1/\gamma)}}{\sqrt{n}}+\frac{(\ell^{2}\kappa^{3}+\ell\kappa^{2})\sqrt{\log(1/\gamma)}}{\sqrt{n}}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\frac{(\ell^{3/2}\kappa^{5/2}+\ell^{1/2}\kappa^{3/2})\log^{1/4}(1/\gamma)}{n^{1/4}}\bigg)~,$
where the alst inequality is by Lemma 8.9 with a domain bound $\|\mathbf{x}_{\mathrm{out}}-\mathbf{x}_{0}\|\leq D$ for some sufficiently large $D$ which is polynomial in all problem parameters (therefore only affecting log terms). Theorem 6.1 follows from applying Theorems 4.1 and 5.1 to bound $\left\|\mathcal{G}_{F_{S},\eta}(\mathbf{x}_{\mathrm{out}})\right\|$ and simplifying.
9 Discussion
In this paper, we studied DP bilevel optimization, and proposed the first algorithms to solve this problem designed for the central DP model, while also being the first that use only gradient queries. Our provided guarantees hold both for constrained and unconstrained settings, cover empirical and population losses alike, and account for mini-batched gradients. As an application, we derived a private update rule for tuning a regularization hyperparameter when fitting a statistical model.
Our work leaves open several directions for future research. First, it is natural to ask what is the extent to which our results can be improved. Notably, even in the well-studied setting of single-level smooth-nonconvex DP optimization, there still exist gaps between known upper and lower bounds for minimizing the gradient norm (cf. Lowy et al. 2024 and discussion therein). The best known lower bound for such problems, which trivially applies also for DP BO which is a strictly more general problem setting, is $\Omega(\sqrt{d}/\epsilon n)$, hinting that our results are perhaps not tight. Moreover, Lowy and Liu (2025) recently presented improved gradient bounds for DP BO via second-order methods, and it is interesting to ask whether first-order methods for this setting, as studied in this work, can be improved.
To follow up on this question, we remark that a candidate strategy to improve the convergence rate in this work would be to use variance reduction, as Arora et al. (2023) used variance reduction to derive faster convergence to private stationarity for single-level DP nonconvex optimization. Applying this for DP BO as the outer loop would require, according to our analysis, to evaluate the cost of inexact gradients in variance-reduced methods, which is left for future work.
Another direction for future work is extending the analysis in Section 7 to derive update rules for privately tuning multiple hyperparameters simultaneously. It is interesting to note that the complexity of our results scales polynomially with the upper-level dimension, which corresponds to the number of hyperparameters, whereas preforming a grid search scales exponentially with the number of hyperparameters. The trade-off is, of course, convergence to local stationarity instead of global optimality.
Lastly, another open direction is understanding whether mini-batch algorithms can avoid the additive $1/b_{\mathrm{out}}$ factor in the unconstrained case $\mathcal{X}=\mathbb{R}^{d_{x}}$. As previously discussed, for constrained problems, even single-level nonconvex algorithms suffer from this batch dependence (Ghadimi et al., 2016). However, for unconstrained problems, Ghadimi and Lan (2013) showed that setting a smaller stepsize, on the order of $\alpha^{2}/\sigma^{2}$, converges to a point with gradient bounded by $\alpha$ after $\mathcal{O}(\alpha^{-4})$ steps, even for $b_{\mathrm{out}}=1$. Applying this to DP bilevel unconstrained optimization seems feasible, and requires accounting for the larger privacy loss due to the slower convergence rate (compared to $\mathcal{O}(\alpha^{-2})$ in our case).
Acknowledgments
This research is supported through an Azrieli Foundation graduate fellowship. We would like to express our gratitude to Vitaly Feldman, Kunal Talwar and the Apple MLR team for the encouraging environment in which this research was initiated. Specifically, we thank Kunal Talwar for insightful discussions, which included pointing out the utility of localizing DP-SGD for optimal high probability guarantees.
Appendix A Differential privacy preliminaries
We recall here some well-known DP facts. The basic composition property of DP states that the (possibly adaptive) composition of $(\epsilon_{0},\delta_{0})$-DP- and $(\epsilon_{1},\delta_{1})$-DP mechanisms, is $(\epsilon_{0}+\epsilon_{1},\delta_{0}+\delta_{1})$-DP. Advanced composition improves this scaling in typical parameter regimes of interest:
Lemma A.1 (Advanced composition, Dwork et al., 2010).
For $\epsilon_{0}<1$, a $T$-fold (possibly adaptive) composition of $(\epsilon_{0},\delta_{0})$-DP mechanisms is $(\epsilon,\delta)$-DP for $\epsilon=\sqrt{2T\log(1/\delta_{0})}\epsilon_{0}+2T\epsilon_{0}^{2},~\delta=(T+1)\delta_{0}$.
We remark that advanced composition is typically used when $\epsilon_{0}\lesssim\sqrt{\log(1/\delta_{0})/T}$, thus the accumulated privacy scales as $\epsilon\asymp\sqrt{T}\epsilon_{0}$.
Lemma A.2 (Gaussian mechanism).
Given a function $h:\Xi^{b}\to\mathbb{R}^{d}$, the Gaussian mechanism $\mathcal{M}(h):\Xi^{b}\to\mathbb{R}^{d}$ defined as $\mathcal{M}(h)(S):=h(S)+\mathcal{N}(\mathbf{0},\sigma^{2}I_{d})$ is $(\epsilon,\delta)$-DP for $\epsilon,\delta\in(0,1)$, as long as $\sigma^{2}\geq\frac{2\log(5/4\delta)(\mathcal{S}_{h})^{2}}{\epsilon^{2}}$, where $\mathcal{S}_{h}:=\sup_{S\sim S^{\prime}}\left\|h(S)-h(S^{\prime})\right\|$ is the $L_{2}$-sensitivity of $h$.
Lemma A.3 (Privacy amplification, Balle et al., 2018).
Suppose $\mathcal{M}:\Xi^{b}\to\mathcal{R}$ is $(\epsilon_{0},\delta_{0})$-DP. Then given $n\geq b$, the mechanism $\mathcal{M}^{\prime}:\Xi^{n}\to\mathcal{R},~\mathcal{M}^{\prime}(S):=\mathcal{M}(B)$ where $B\sim\mathrm{Unif}(\Xi)^{b}$, is $(\epsilon,\delta)$-DP for $\epsilon=\log(1+(1-(1-1/n)^{b})(e^{\epsilon_{0}}-1)),~\delta=\delta_{0}$.
Privacy amplification is typically used when $\epsilon_{0}\leq 1$, under which the privacy after subsampling scales as $\epsilon\asymp\frac{b\epsilon_{0}}{n}$ (since $e^{\epsilon_{0}}-1\asymp\epsilon_{0},~(1-1/n)^{b}\asymp\frac{b}{n}$ and $\log(1+\frac{b}{n}\epsilon_{0})\asymp\frac{b}{n}\epsilon_{0}$).
Appendix B Optimal DP algorithm for strongly-convex objectives
The goal of this appendix is to provide a self contained analysis of a DP algorithm for strongly-convex optimization which achieves the optimal convergence rate with a a high probability guarantee. Any such algorithm can be used as the inner loop in our DP bilevel algorithm.
In particular, we analyze localized DP (S)GD. Although it would have been more natural to apply DP-SGD, this seems (at least according to our analysis) to yield an inferior rate with respect to the required high probability guarantee. [^7] Even if we would have sought only expectation bounds with respect to the hyperobjective, the high probability bound with respect to the inner problem is key to being able to argue about the gradient inexactness thereafter. Indeed, for DP-(S)GD, previous works (such as Bassily et al. 2014) typically provide bounds in expectation, and then convert them into high-probability bounds via a black-box reduction, which applies several runs and selects the best run via the private noisy-min (via Laplace mechanism). The additional error incurred by this selection is of order $\frac{1}{n}$, which translates to $\frac{1}{\sqrt{n}}$ in terms of distance to the optimum, thus spoiling the fast rate of $\frac{1}{n}$ otherwise achieved in expectation for strongly-convex objectives. We therefore resort to localization [Feldman et al., 2020]: by running projected-(S)GD over balls with shrinking radii, applying martingale concentration bounds enables us to show that the distance to optimum shrinks as $R_{m+1}\lesssim\sqrt{\frac{R_{m}}{n}}+\frac{1}{n}$, and thus with negligible overhead we eventually recover the optimal fast rate $R_{M}\lesssim\frac{1}{n}$ with high probability. Our analysis differs than previous localization analyses, as it does not require adapting the noise-level and step sizes throughout the rounds.
We prove the following (which easily implies also the full-batch Theorem 8.2):
Theorem B.1.
Suppose that $h:\mathbb{R}^{d_{y}}\times\Xi\to\mathbb{R}$ is a $\mu$-strongly-convex function of the form $h(\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}h(\mathbf{y},\xi_{i})$ where $h(\cdot,\xi_{i})$ is $L$-Lipschitz for all $i\in[n]$. Suppose $\arg\min h=:\mathbf{y}^{*}\in\mathbb{B}(\mathbf{y}_{0},R_{0})$, and that $n\geq\frac{LR_{0}^{\frac{2}{\log(d_{y})}}}{\mu\epsilon^{\prime}}$. Then given any batch size $b\in\{1,\dots,n\}$, there is an assignment of parameters $M=\log_{2}\log(\frac{\mu\epsilon^{\prime}n}{L}),~\sigma^{2}_{\mathrm{SGD}}=\widetilde{\mathcal{O}}\left(\frac{L^{2}}{\epsilon^{\prime 2}}\right),~\eta_{t}=\frac{1}{\mu(t+1)},~T_{\mathrm{SGD}}=n^{2},~R_{m}=\widetilde{\Theta}\left(\sqrt{\frac{R_{m-1}L}{\mu\epsilon^{\prime}n}}+\frac{L\sqrt{d_{y}}}{\mu\epsilon^{\prime}n}\right)$ such that running Algorithm 4 satisfies $(\epsilon,\delta)$-DP, and outputs $\mathbf{y}_{\mathrm{out}}$ such that $\left\|\mathbf{y}_{\mathrm{out}}-\mathbf{y}^{*}\right\|=\widetilde{\mathcal{O}}\left(\frac{L\sqrt{d_{y}}}{\mu n\epsilon}\right)$ with probability at least $1-\gamma$.
Proof of Theorem B.1.
[…]
We turn to prove the utility of the algorithm. We first show that for all $m:$
$\Pr\left[\mathbf{y}^{*}\in\mathbb{B}(\mathbf{y}_{0}^{m},R_{m})\right]\geq 1-\frac{m\gamma}{M}~.$ (18)
We prove this by induction over $m$. The base case $m=0$ follows by the assumption $\mathbf{y}^{*}\in\mathbb{B}(\mathbf{y}_{0},R_{0})$. Denoting $\mathbf{e}_{t}^{m}:=\nabla h(\mathbf{y}_{t}^{m};B_{t})-\nabla h(\mathbf{y}_{t}^{m})$, using the inductive hypothesis that $\mathbf{y}^{*}\in\mathbb{B}(\mathbf{y}_{0}^{m},R_{m})$ with probability at least $1-\frac{m\gamma}{M}$, under this probably event we get
$\left\|\mathbf{y}_{t+1}^{m}-\mathbf{y}^{*}\right\|^{2}$ $=\left\|\mathrm{Proj}_{\mathbb{B}(\mathbf{y}_{0}^{m},R_{m})}\left[\mathbf{y}_{t}^{m}-\eta_{t}(\nabla h(\mathbf{y}_{t}^{m};B_{t}^{m})+\nu_{t}^{m})\right]-\mathbf{y}^{*}\right\|^{2}$
$\leq\left\|\mathbf{y}_{t}^{m}-\eta_{t}(\nabla h(\mathbf{y}_{t}^{m};B_{t}^{m})+\nu_{t}^{m})-\mathbf{y}^{*}\right\|^{2}$
$=\left\|\mathbf{y}_{t}^{m}-\mathbf{y}^{*}\right\|^{2}-2\eta_{t}\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\nabla h(\mathbf{y}_{t}^{m};B_{t}^{m})+\nu_{t}^{m}\right\rangle+\eta_{t}^{2}\left\|\nabla h(\mathbf{y}_{t}^{m};B_{t}^{m})+\nu_{t}^{m}\right\|^{2}$
$\leq\left\|\mathbf{y}_{t}^{m}-\mathbf{y}^{*}\right\|^{2}-2\eta_{t}\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\nabla h(\mathbf{y}_{t})+\mathbf{e}^{m}_{t}+\nu_{t}^{m}\right\rangle+2\eta_{t}^{2}\left(\left\|\nu^{m}_{t}\right\|^{2}+\left\|\nabla h(\mathbf{y}_{t})\right\|^{2}\right)$
$=\left\|\mathbf{y}_{t}^{m}-\mathbf{y}^{*}\right\|^{2}-2\eta_{t}\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\nabla h(\mathbf{y}_{t}^{m})\right\rangle$
$~~~~~-2\eta_{t}\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\mathbf{e}^{m}_{t}+\nu_{t}^{m}\right\rangle+2\eta_{t}^{2}\left(\left\|\nu^{m}_{t}\right\|^{2}+\left\|\nabla h(\mathbf{y}_{t})\right\|^{2}\right)~.$
Rearranging, and using the strong convexity and Lipschitz assumptions, we see that
$h(\mathbf{y}_{t}^{m})-h(\mathbf{y}^{*})$ $\leq\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\nabla h(\mathbf{y}_{t}^{m})\right\rangle-\frac{\mu}{2}\left\|\mathbf{y}_{t}^{m}-\mathbf{y}^{*}\right\|^{2}$
$\leq\left(\frac{1}{2\eta_{t}}-\frac{\mu}{2}\right)\left\|\mathbf{y}_{t}^{m}-\mathbf{y}^{*}\right\|^{2}-\frac{1}{2\eta_{t}}\left\|\mathbf{y}_{t+1}^{m}-\mathbf{y}^{*}\right\|^{2}$
$~~~~~~-\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\mathbf{e}^{m}_{t}+\nu_{t}^{m}\right\rangle+\eta_{t}\left(\left\|\nu_{t}^{m}\right\|^{2}+L^{2}\right)~.$
Averaging over $t$ and using $\eta_{t}=\frac{1}{\mu(t+1)}$, which satisfies $\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}-\mu\right)\leq 0$ and $\frac{1}{T}\sum_{t=0}^{T}\eta_{t}\lesssim\frac{\log T}{\mu T}$, by Jensen’s inequality, overall we get with probability at least $1-\frac{m\gamma}{M}:$
$h(\mathbf{y}_{0}^{m+1})-h(\mathbf{y}^{*})$ $=h\left(\frac{1}{T}\sum_{t=0}^{T-1}\mathbf{y}_{t}^{m}\right)-h(\mathbf{y}^{*})$
$\lesssim\underset{(I)}{\underbrace{\left|\frac{1}{T}\sum_{t=0}^{T-1}\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\mathbf{e}^{m}_{t}+\nu_{t}^{m}\right\rangle\right|}}+\frac{L^{2}\log T}{\mu T}+\frac{\log T}{\mu T}\underset{(II)}{\underbrace{\sum_{t=0}^{T-1}\left\|\nu_{t}^{m}\right\|^{2}}}~$. (19)
We now apply concentration inequalities to bound $(I)$ and $(II)$ with high probability, for which we will use basic properties of sub-Gaussian distributions (cf. Vershynin 2018, §3.4). To bound $(I)$, note that for all $t:~\mathbb{E}\mathbf{e}^{m}_{t}=\mathbb{E}\nu^{m}_{t}=\mathbf{0}$ and therefore $\mathbb{E}\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\mathbf{e}^{m}_{t}+\nu_{t}^{m}\right\rangle=0$. Moreover, $\mathbf{e}^{m}_{t}=\frac{1}{b}\sum_{\xi\in B_{t}^{m}}(\nabla h(\mathbf{x}_{t}^{m};\xi)-\nabla h(\mathbf{y}_{t}^{m}))$ is the average of $b$ independent vectors with norm bounded by at most $2L$, while $\nu_{t}^{m}\sim\mathcal{N}(\mathbf{0},\sigma^{2}_{\mathrm{SGD}}I_{d_{y}})=\mathcal{N}(\mathbf{0},\widetilde{\mathcal{O}}(\frac{L^{2}}{\epsilon^{\prime 2}})\cdot I_{d_{y}})$, and also $\left\|\mathbf{y}_{t}^{m}-\mathbf{y}^{*}\right\|\leq R_{m}$ by the inductive hypothesis. By combining all of these observations, we see that $\left\langle\mathbf{y}_{t}^{m}-\mathbf{y}^{*},\mathbf{e}^{m}_{t}+\nu_{t}^{m}\right\rangle$ is a $\mathcal{O}(R_{m}\cdot(\frac{L}{b}+\frac{L}{\epsilon^{\prime}}))=\mathcal{O}(\frac{R_{m}L}{\epsilon^{\prime}})$-sub-Gaussian random variable. By Azuma’s inequality for sub-Gaussians [Shamir, 2011], we get that with probability at least $1-\frac{\gamma}{2M}:$
$(I)=\widetilde{\mathcal{O}}\left(\frac{\frac{R_{m}L}{\epsilon^{\prime}}\log(\gamma/M)}{\sqrt{T}}\right)=\widetilde{\mathcal{O}}\left(\frac{R_{m}L}{\epsilon^{\prime}n}\right)~.$ (20)
To bound $(II)$, by concentration of the Gaussian norm (cf. Vershynin 2018, Theorem 3.1.1) and the union bound we can get that with probability at least $1-\frac{\gamma}{2M}:$
$(II)=\widetilde{\mathcal{O}}\left(d_{y}\sigma_{\mathrm{SGD}}^{2}\right)=\widetilde{\mathcal{O}}\left(\frac{d_{y}L^{2}}{\epsilon^{\prime 2}}\right)~.$ (21)
Plugging Eqs. (20) and (21) into Eq. (19), we overall get that with probability at least $1-\frac{m\gamma}{M}-2\cdot\frac{\gamma}{2M}=1-\frac{(m+1)\gamma}{M}:$
$h(\mathbf{y}_{0}^{m+1})-h(\mathbf{y}^{*})=\widetilde{\mathcal{O}}\left(\frac{R_{m}L}{\epsilon^{\prime}n}+\frac{d_{y}L^{2}}{\mu\epsilon^{\prime 2}n^{2}}\right)~.$
Applying the $\mu$-strong-convexity of $h$, and sub-additivity of the square root, we get that
$\left\|\mathbf{y}_{0}^{m+1}-\mathbf{y}^{*}\right\|\leq\sqrt{\frac{2(h(\mathbf{y}_{0}^{m+1})-h(\mathbf{y}^{*}))}{\mu}}=\widetilde{\mathcal{O}}\left(\sqrt{\frac{R_{m}L}{\mu\epsilon^{\prime}n}}+\frac{L\sqrt{d_{y}}}{\mu\epsilon^{\prime}n}\right)\leq R_{m+1}~.$
We have therefore proven Eq. (18). In particular for $m=M$ we get that with probability at least $1-\gamma:$
$\left\|\mathbf{y}_{\mathrm{out}}-\mathbf{y}^{*}\right\|\leq R_{M}~,$ (22)
hence it remains to bound $R_{M}$. We will prove, once again by induction over $m$, that
$R_{m}=\widetilde{\mathcal{O}}\left(R_{0}^{\frac{1}{2^{m}}}\left(\frac{L}{\mu\epsilon^{\prime}n}\right)^{1-\frac{1}{2^{m}}}+\frac{L}{\mu\epsilon^{\prime}n}\sum_{i=1}^{m}d_{y}^{\frac{1}{2^{i}}}\right)~.$ (23)
The base $m=0$ simply follows since the left hand side in Eq. (23) reduces to $R_{0}$. Denoting $A:=\frac{L}{\mu\epsilon^{\prime}n}$, by our assignment of $R_{m+1}$, the induction hypothesis and sub-additivitiy of the square root we get:
$R_{m+1}$ $=\widetilde{\mathcal{O}}\left(\sqrt{R_{m}A}+A\sqrt{d_{y}}\right)$
$=\widetilde{\mathcal{O}}\left(A^{1/2}\left(R_{0}^{\frac{1}{2^{m+1}}}A^{\frac{1}{2}-\frac{1}{2^{m+1}}}+A^{1/2}\sum_{i=1}^{m}d_{y}^{\frac{1}{2^{i+1}}}\right)+Ad_{y}^{1/2}\right)$
$=\widetilde{\mathcal{O}}\left(R_{0}^{\frac{1}{2^{m+1}}}A^{1-\frac{1}{2^{m+1}}}+A\sum_{i=1}^{m+1}d_{y}^{\frac{1}{2^{i}}}\right)~,$
therefore proving Eq. (23). In particular, for $m=M=\log_{2}\log(\frac{\mu\epsilon^{\prime}n}{L})$, which satisfies $\frac{1}{2^{M}}=\frac{1}{\log(\frac{\mu\epsilon^{\prime}n}{L})}=\frac{1}{\log(1/A)}$ we get
$R_{M}$ $=\widetilde{\mathcal{O}}\left(R_{0}^{\frac{1}{\log(1/A)}}A^{1+\frac{1}{\log(A)}}+MAd_{y}^{1/2}\right)$
$=\widetilde{\mathcal{O}}\left(R_{0}^{\frac{1}{\log(1/A)}}A+Ad^{1/2}\right)$
$=\widetilde{\mathcal{O}}\left(R_{0}^{\frac{1}{\log(\mu\epsilon^{\prime}n/L)}}\frac{L}{\mu\epsilon^{\prime}n}+\frac{Ld_{y}^{1/2}}{\mu\epsilon^{\prime}n}\right)$
$=\widetilde{\mathcal{O}}\left(\frac{L\sqrt{d_{y}}}{\mu\epsilon^{\prime}n}\right)~,$
where the last follows from our assumption on $n$. This completes the proof by Eq. (22). ∎
Appendix C Auxiliary lemma
We will recall a useful fact, which asserts that the mapping $\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x})$ is non-expansive with respect to $\mathbf{v}:$
Lemma C.1.
For any $\mathbf{x},\mathbf{v},\mathbf{w}\in\mathbb{R}^{d},~\eta>0:~\left\|\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{G}_{\mathbf{w},\eta}(\mathbf{x})\right\|\leq\left\|\mathbf{v}-\mathbf{w}\right\|$.
In our analysis we argue that in the bilevel setting, gradient estimates must be inexact to avoid privacy leaking between the two levels, and Lemma C.1 allows us to control the error due to this inexactness. Lemma C.1 is known (cf. Ghadimi et al. 2016), and we reprove it here for completeness.
Proof of Lemma C.1.
The proof is due to Ghadimi et al. [2016]. By definition,
$\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})$ $=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\langle\mathbf{v},\mathbf{u}\rangle+\frac{1}{2\eta}\left\|\mathbf{x}-\mathbf{u}\right\|^{2}\right\}~,$
$\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})$ $=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\langle\mathbf{w},\mathbf{u}\rangle+\frac{1}{2\eta}\left\|\mathbf{x}-\mathbf{u}\right\|^{2}\right\}~,$
hence by first order optimality criteria, for any $\mathbf{u}\in\mathcal{X}:$
$\left\langle\mathbf{v}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathbf{x}),\mathbf{u}-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle$ $\geq 0~,$
$\left\langle\mathbf{w}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathbf{x}),\mathbf{u}-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})\right\rangle$ $\geq 0~.$
Plugging $\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})$ as $\mathbf{u}$ into the first inequality above, and $\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})$ into the second, shows that
$0$ $\leq\left\langle\mathbf{v}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathbf{x}),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle~,$
$0$ $\leq\left\langle\mathbf{w}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathbf{x}),\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})\right\rangle=\left\langle-\mathbf{w}+\frac{1}{\eta}(\mathbf{x}-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle~.$
Summing the two inequalities yields
$0$ $\leq\left\langle{\mathbf{v}-\mathbf{w}}+\frac{1}{\eta}(\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})),\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle$
$=\left\langle\mathbf{v}-\mathbf{w},\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\rangle-\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|^{2}$
$\leq\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|\left(\left\|\mathbf{v}-\mathbf{w}\right\|-\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})\right\|\right)~.$
Hence,
$\left\|\mathbf{v}-\mathbf{w}\right\|$ $\geq\frac{1}{\eta}\left\|\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})\right\|$
$=\left\|\frac{1}{\eta}\left(\mathcal{P}_{\mathbf{v},\eta}(\mathbf{x})-\mathbf{x}\right)-\frac{1}{\eta}\left(\mathcal{P}_{\mathbf{w},\eta}(\mathbf{x})-\mathbf{x}\right)\right\|$
$=\left\|\mathcal{G}_{\mathbf{v},\eta}(\mathbf{x})-\mathcal{G}_{\mathbf{w},\eta}(\mathbf{x})\right\|~.$